### v2

# Matroid Products via Submodular Coupling

Abstract
The study of matroid products traces back to the 1970s, when Lovász and Mason studied the existence of various types of matroid products with different strengths. Among these, the tensor product is arguably the most important, which can be considered as an extension of the tensor product from linear algebra. However, Las Vergnas showed that the tensor product of two matroids does not always exist. Over the following four decades, matroid products remained surprisingly underexplored, regaining attention only in recent years due to applications in tropical geometry and the limit theory of matroids.
In this paper, inspired by the concept of coupling in probability theory, we introduce the notion of coupling for matroids – or, more generally, for submodular set functions. This operation can be viewed as a relaxation of the tensor product. Unlike the tensor product, however, we prove that a coupling always exists for any two submodular functions and can be chosen to be increasing if the original functions are increasing. As a corollary, we show that two matroids always admit a matroid coupling, leading to a novel operation on matroids. Our construction is algorithmic, providing an oracle for the coupling matroid through a polynomial number of oracle calls to the original matroids.
We apply this construction to derive new necessary conditions for matroid representability and establish connection between tensor products and Ingleton’s inequality. Additionally, we verify the existence of set functions that are universal with respect to a given property, meaning any set function over a finite domain with that property can be obtained as a quotient.
Keywords: Amalgams, Coupling, Coverage functions, Matroids, Product, Quasi-product, Quotients, Submodular functions, Tensor product
 [^1] An extended abstract [6] of the paper appeared at 57th Annual ACM Symposium on Theory of Computing (STOC 2025).
1 Introduction
Coupling of probability measures, generally attributed to Doeblin [11], is a key concept in probability theory through which random variables can be compared with each other. For two probability measures, a coupling is a joint probability measure over the product of the underlying spaces, such that the marginals correspond to the given probability measures. Formally, if $\mu_{1}$ and $\mu_{2}$ are probability measures on spaces $(S_{1},\mathcal{B}_{1})$ and $(S_{2},\mathcal{B}_{2})$, respectively, a coupling is a joint probability measure $\mu$ on the product space $(S_{1}\times S_{2},\mathcal{B}_{1}\otimes\mathcal{B}_{2})$ such that the marginal of $\mu$ on $S_{i}$ is $\mu_{i}$ for $i=1,2$, i.e., $\mu(X_{1}\times S_{2})=\mu_{1}(X_{1})$ for any $X_{1}\in\mathcal{B}_{1}$ and $\mu(S_{1}\times X_{2})=\mu_{2}(X_{2})$ for any $X_{2}\in\mathcal{B}_{2}$. Usually, the goal is to choose a coupling that allows for comparisons between the probability measures, such as minimizing differences or distances between them. Coupling techniques have found applications in various areas. In optimal transport theory, the goal is to find the most efficient way to transform one distribution into another, with applications in economics, statistical physics and machine learning [9]. In stochastic processes, coupling helps to describe the joint behavior of random variables and to analyze and bound convergence rates in chains with more complex dependencies [27, 1].
Interestingly, a similar notion has also appeared in a combinatorial line of research on matroids. Let $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$ be finite matroids over ground sets $S_{1}$ and $S_{2}$ with rank functions $r_{1}$ and $r_{2}$, respectively. A matroid $M=(S_{1}\times S_{2},r)$ is called a quasi-product of $M_{1}$ and $M_{2}$ if the restriction of $M$ to $\{x\}\times S_{2}$ is isomorphic to $M_{2}$ by the natural bijection with $S_{2}$ for all non-loops $x\in S_{1}$ and is the zero matroid for loops $x$, and analogously with the two factors interchanged. If furthermore $X\times Y$ is a flat of $M$ for all flats $X$ of $M_{1}$ and $Y$ of $M_{2}$, then $M$ is called a product of $M_{1}$ and $M_{2}$. Finally, if $r(X\times Y)=r_{1}(X)\cdot r_{2}(Y)$ holds for every $X\subseteq S_{1},Y\subseteq S_{2}$, then $M$ is a tensor product of $M_{1}$ and $M_{2}$. Note that at this point, it is not clear whether any two matroids admit a tensor product, a product, or even a quasi-product. Mason [32] and Lovász [28] provided constructions yielding a product of rank $r_{1}(S_{1})+r_{2}(S_{2})-1$ via Dilworth completion. The existence of a tensor product, however, was asked as an open question in [28]. Note that if both $M_{1}$ and $M_{2}$ are linear matroids over the same field, then it is easy to construct a tensor product by taking the tensor product of the matrices representing them. Nevertheless, Las Vergnas [26] later showed that the situation is generally not so fortunate, as, for example, the Vámos matroid and the rank-2 uniform matroid on three elements do not admit a tensor product.
One can think of the tensor product of two matroids as a construction where the product of independent sets is itself independent. However, since a tensor product does not exist for every pair of matroids, Mason [31] raised the question of whether a freest product exists among the possible products. Las Vergnas [26] answered this question in the negative and showed that the matroid obtained via Dilworth truncation as in [28, 31] is the freest quasi-product among those satisfying $r(\{x,x^{\prime}\}\times\{y,y^{\prime}\})\leq 3$ for all $x,x^{\prime}\in S_{1}$ and $y,y^{\prime}\in S_{2}$, where $r$ denotes the rank function of the product matroid.
1.1 Our results and techniques
In this paper, we extend the concept of coupling to matroids and, more broadly, to submodular functions. We prove that, unlike tensor products, couplings always exist, leading to a novel operation on matroids. This result is particularly interesting because it shows that a finite number of matroids can be encoded into a single one, while also taking into account how these matroids interact with each other. As an application, we establish the existence of functions that are universal with respect to some property, meaning that any set function with that property can be obtained as a quotient. This result brings new insights into the developing limit theory of matroids and submodular functions.
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ be set functions defined over finite ground sets $S_{1}$ and $S_{2}$, respectively. The notion of a tensor product can be naturally extended to set functions by calling a function $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ a tensor product of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times X_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(X_{2})$ holds for every $X_{1}\subseteq S_{1}$ and $X_{2}\subseteq S_{2}$. As a new concept, we call $\varphi$ a coupling of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times S_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(S_{2})$ for every $X_{1}\subseteq S_{1}$ and $\varphi(S_{1}\times X_{2})=\varphi_{1}(S_{1})\cdot\varphi_{2}(X_{2})$ for every $X_{2}\subseteq S_{2}$. In other words, a coupling is a set function on the product set, where the projections onto each coordinate return the corresponding $\varphi_{i}$, up to a constant multiplier. It is worth emphasizing that neither a tensor product nor a coupling is uniquely determined from the factors $\varphi_{1}$ and $\varphi_{2}$.
Motivated by the goal of defining the coupling of matroids, Section 3 focuses on submodular functions. First, we show that any two submodular functions admit a submodular coupling (Theorem 3.1). One remarkable feature of the proof is that it provides an explicit formula for the coupling function using two arbitrary modular functions. However, the proof itself does not imply that if both $\varphi_{1}$ and $\varphi_{2}$ are monotonically increasing, then the coupling function $\varphi$ is also monotonically increasing. Since this property is essential for the coupling of two matroids to result in a matroid, next we show that a $k_{1}$-polymatroid function and a $k_{2}$-polymatroid function have a $(k_{1}\cdot k_{2})$-polymatroid coupling, which is also integer-valued if the original functions are (Theorem 3.2). The proof follows a similar idea as that of the submodular case, but the modular functions in question need to be chosen to be elements in the base polyhedra of $\varphi_{1}$ and $\varphi_{2}$, respectively. As a corollary, we get that any two matroids admit a coupling (Corollary 3.6) which in turn implies that there exists a finite matroid that contains every matroid of fixed rank over a ground set of fixed size (Corollary 3.7). We provide several equivalent characterizations of two matroids having a coupling that is also a tensor product (Theorem 3.11), and establish a strong connection between tensor products and Ingleton’s inequality, a fundamental tool in the theory of representable matroids (Theorem 3.14). As a corollary, we get a new proof for the fact that the uniform matroid $U_{2,3}$ and the Vámos matroid do not admit a tensor product. We separately study a subclass of increasing submodular functions with strong structural properties, called “coverage functions” in combinatorial optimization. These functions were originally introduced by Choquet for analytical studies. He defined them in terms of a sequence of inequalities strengthening submodularity, and proved their equivalence with what is now the combinatorial definition. We prove that a coverage function and an increasing submodular function have an increasing submodular tensor product (Corollary 3.17), and that any two coverage functions have a tensor product that is a coverage function (Corollary 3.19). The proofs rely on the characterization of the extreme rays of the cone of coverage functions, as established by Choquet. Finally, we describe how to extend our results on submodular functions defined on finite ground sets to those with infinite domains (Theorems 3.22 and  3.23).
Let $\varphi$ be a set function over some finite ground set $S$, and let $\mathcal{Q}=(S_{1},\dots,S_{q})$ be a partition of $S$ into $q$ possibly empty parts. The quotient of $\varphi$ with respect to $\mathcal{Q}$ is a set function $\varphi_{\mathcal{Q}}$ over $S_{\mathcal{Q}}=\{s_{1},\dots,s_{q}\}$ defined by $\varphi_{\mathcal{Q}}(X)=\varphi(\bigcup_{s_{i}\in X}S_{i})$ for $X\subseteq S_{\mathcal{Q}}$. In particular, if $\varphi\colon 2^{S_{1}\times S_{2}}\to\mathbb{R}$ is a coupling of $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}$ and $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are the partitions of $S_{1}\times S_{2}$ into fibers of the form $\{x\}\times S_{2}$ and $S_{1}\times\{y\}$, respectively, then $\varphi_{1}=\varphi_{\mathcal{Q}_{1}}/\varphi_{2}(S_{2})$ and $\varphi_{2}=\varphi_{\mathcal{Q}_{2}}/\varphi_{1}(S_{1})$. The positive results of Section 3 motivate the following question: Given a property of set functions, is there a function that is universal with respect to that property, in the sense that every function exhibiting the property can be obtained as its quotient?
In Section 4, we examine this problem and show that, under some natural assumptions, the answer is positive (Theorem 4.2). Although the proof is analytical and may therefore be less appealing to those interested in combinatorial optimization, the result has far-reaching implications for functions defined on finite ground sets as well. For example, we obtain that there exists a submodular function $\varphi\colon 2^{\mathbb{Z}}\to[0,1]$ such that any nonnegative normalized submodular function is a quotient of $\varphi$ (Corollary 4.3), and we prove an analogous result for coverage functions as well (Corollary 4.4). We conclude the section by providing an explicit construction for a universal coverage function, being in sharp contrast with the preceding non-constructive universality results; the formula of the function is surprisingly simple (Theorem 4.6).
It is worth emphasizing that for functions over finite ground sets, our proofs are algorithmic in the sense that given value oracles for $\varphi_{1}$ and $\varphi_{2}$, one can construct a value oracle for their coupling $\varphi$ with the desired properties in polynomial time.
1.2 Related work and motivation
Understanding how matroids can be combined does not only help in studying properties of larger systems formed from simpler matroid components, but it is closely related to a range of problems and techniques in combinatorial optimization. In what follows, we give a brief overview of related topics.
Tropical geometry.
Tensor products of matroids have deep connections to tropical geometry. A key consequence of Las Vergnas’s counterexample is that the tensor product of tropical linear spaces does not always yield a tropical linear space, see e.g. [19]. The notion of tensor products can be extended to more than two matroids. If all $k$ matroids in the product are identical to a given matroid $M=(S,r)$, the tensor product defines a matroid structure on the set of ordered $k$-tuples from $S$, known as the $k$-th power of $M$. For unordered $k$-tuples, we consider symmetric tensors, which remain invariant under any permutation of their variables. The concept of symmetric powers of matroids was first introduced by Lovász [28] and Mason [31], who noted that not all matroids admit higher symmetric powers. Draisma and Rincón [14] established a link between tropical ideals and matroid symmetric powers by showing that the Bergman fan of the direct sum of the Vámos matroid and $U_{2,3}$ is not a tropically realizable variety. In [2], Anderson proved an equivalence between valuated matroids with arbitrarily large symmetric powers and tropical linear spaces represented as varieties of tropical ideals. Brakensiek, Dhar, Gao, Gopi, and Larson [8] explored the connection between rigidity matroids of graphs and matroids arising from linear algebraic constructions like tensor products and symmetric products. Matroid tensor products in tropical geometry were also examined in [23, 15].
Linear matroids.
A matroid is called linear if its independent sets can be represented by vectors in a vector space over some field $\mathbb{F}$. Such matroids are particularly interesting because they allow for a rich connection between linear algebra and combinatorial structures. For deciding if a matroid is regular, i.e. representable over every field, Seymour’s decomposition theorem [36] implies an algorithm, as such decompositions can be found efficiently [37]. However, Truemper [37] showed that many representability questions cannot be efficiently solved. This includes deciding representability over a specific field, over all fields with a given characteristic, and, most importantly, over any field. Since determining representability is oracle-hard in general, the focus of research has shifted toward finding necessary or sufficient conditions. Ingleton’s inequality [22] is one of the most prominent examples, giving a necessary condition for a matroid to be linear. Surprisingly, we show that representability of a matroid is closely related to having a tensor product with $U_{2,3}$.
Limits of matroids.
The limit theory of graphs provides powerful tools for analyzing sequences of graphs and their structural similarities through analytic methods. By defining convergence in terms of distributions of small subgraphs (left-convergence) and homomorphisms into small graphs (right-convergence), it aids in understanding complex networks and their applications in various fields of mathematics and computer science; we refer the interested reader to [29] for a thorough introduction. In [5], a new form of right-convergence called quotient-convergence was introduced for set functions, which eventually led to a notion of convergence of matroids through their rank functions. The limit object of such a sequence is a submodular function [30]. One of the main research subjects is understanding the structure and properties of these limit objects, which heavily relies on examining their quotients and the coupling of the functions in the sequence.
Extended formulations.
For a polytope $P\subseteq\mathbb{R}^{d}$, an extension of $P$ is a polytope $Q\in\mathbb{R}^{d^{\prime}}$ such that $P=\pi(Q)$ for some affine map $\pi\colon\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d}$. The extension complexity $\operatorname{xc}(P)$ of $P$ is defined as the minimum number of facets of an extension of $P$. If $Q$ is given by a linear description $Q=\{y\in\mathbb{R}^{d^{\prime}}\mid Ay\leq b\}$, then $\{x\in\mathbb{R}^{d}\mid Ay\leq b,x=\pi(y)\}$ is called an extended formulation of $P$. Using this terminology, the extension complexity of a polytope is the minimum number of inequality constraints in an extended formulation. Since the fundamental work of Yannakakis [39], the extension complexity of various families of polytopes that appear naturally in combinatorial and graph optimization problems has been settled, such as the traveling salesman and cut polytopes [16], the stable set polytope [16, 20], the matching polytope [35], and recently the independence polytope of regular matroids [3]. Taking the quotient of a function is similar to projection; however, instead of retaining only certain coordinates of a solution, the coordinates are partitioned and summed within each partition class. Consequently, the coupling of matroids, or more generally, submodular functions leads to a different type of “extended” formulation that is of independent combinatorial interest.
1.3 Organization
The rest of the paper is organized as follows. In Section 2, we introduce basic definitions, notation, and relevant results on submodular functions, matroids, and polymatroids. Section 3 is devoted to verifying the existence of couplings for various functions, first focusing on submodular functions in Section 3.1 and then extending these results to polymatroid functions in Section 3.2. In Section 3.3, we show how these observations lead to one of the main results of the paper: the existence of matroidal couplings that are almost tensor products. Section 3.3.1 explores the relation between our construction and matroidal amalgams. To understand the fine line between couplings and tensor products, we establish a necessary and sufficient condition for a coupling to be a tensor product in Section 3.3.2. Finally, in Section 3.3.3, we provide new necessary conditions for matroid representability. As an application, we verify the existence of functions that are universal with respect to certain properties in Section 4. The proof is analytical and hence is not constructive; we encourage first-time readers to skip the technical parts of Sections 4.1 and 4.2. As an application of the abstract existence theorem, we deduce that universal functions exist for several set function properties; these examples can be found in Section 4.3. Finally, we give an explicit construction for a universal coverage function in Section 4.4.
2 Preliminaries
Basic notation.
We denote the sets of reals and integers by $\mathbb{R}$ and $\mathbb{Z}$, and add $+$ or $>0$ as a subscript when considering nonnegatives or strictly positive values only. The cardinality of $\mathbb{R}$ is denoted by $\mathfrak{c}$. For a real number $a\in\mathbb{R}$, we denote its fractional part by $\operatorname{frac}(a)$. For a positive integer $k$, we use $[k]\coloneqq\{1,\dots,k\}$ while $[0]=\emptyset$ by convention. Given a ground set $S$, a subset $X\subseteq S$ and $y\in S$, the sets $X\setminus\{y\}$ and $X\cup\{y\}$ are abbreviated as $X-y$ and $X+y$, respectively. Moreover, we often denote a single element set $\{x\}$ by $x$ when this causes no confusion. The complement of $X$ is denoted by $X^{c}=S\setminus X$. Given a function $\mu\colon S\to\mathbb{R}_{+}$, we use the notation $\mu(X)\coloneqq\sum_{x\in X}\mu(x)$. If $S=S_{1}\times S_{2}$ for some sets $S_{1}$ and $S_{2}$, then for any $Z\subseteq S$, $x\in S_{1}$ and $y\in S_{2}$, we use $Z_{x}=Z\cap(\{x\}\times S_{2})$ and $Z^{y}=Z\cap(S_{1}\times\{y\})$ to denote the $x$- and $y$-fibers of $Z$, respectively. Furthermore, we use $\pi_{i}\colon S\to S_{i}$ for denoting the coordinate maps, that is, $\pi_{1}(Z)=\{x\in S_{1}\mid Z_{x}\neq\emptyset\}$ and $\pi_{2}(Z)=\{y\in S_{2}\mid Z^{y}\neq\emptyset\}$.
Matroids.
For basic definitions on matroids, we refer the reader to [33]. A matroid $M=(S,r)$ is defined by its finite ground set $S$ and its rank function $r\colon 2^{S}\to\mathbb{Z}_{+}$ that satisfies the rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Here, (R1) is a standard normalizing assumption, while (R2) requires the rank function to be increasing, (R3) its subcardinality, and (R4) its submodularity. The rank of the matroid is $r(S)$, and by the normalized rank function of $M$ we mean the set function $r/r(S)$. A subset $X\subseteq S$ is called independent if $|X|=r(X)$. An inclusionwise minimal non-independent set forms a circuit, while a loop is a circuit consisting of a single element. A flat is a set $F\subseteq S$ such that $r(F+e)=r(F)+1$ holds for every $e\in S\setminus F$. A matroid is called modular if $r(F_{1})+r(F_{2})=r(F_{1}\cap F_{2})+r(F_{1}\cup F_{2})$ for any pair of flats $F_{1},F_{2}$. For sets $X,Y\subseteq S$, we say that $X$ spans $Y$ in $M$ if $r(X\cup Y)=r(X)$.
Given a subset $S^{\prime}\subseteq S$, the restriction of $M$ to $S^{\prime}$ is a matroid $M|S^{\prime}=(S^{\prime},r^{\prime})$ whose rank function is the restriction of $r$ to subsets of $S^{\prime}$. The contraction of $S^{\prime}$ results in a matroid $M/S^{\prime}=(S\setminus S^{\prime},r^{\prime})$ with rank function $r^{\prime}(X)=r(X\cup S^{\prime})-r(S^{\prime})$ for each $X\subseteq S\setminus S^{\prime}$. A matroid $N$ that can be obtained from $M$ by a sequence of restrictions and contractions is called a minor of $M$. The direct sum $M_{1}\oplus M_{2}$ of matroids $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$ on disjoint ground sets is the matroid $M=(S_{1}\cup S_{2},r)$ whose independent sets are the disjoint unions of an independent set of $M_{1}$ and an independent set of $M_{2}$, that is, $r(X)=r_{1}(X\cap S_{1})+r_{2}(X\cap S_{2})$ for $X\subseteq S_{1}\cup S_{2}$.
A matroid is linear or representable over some field $\mathbb{F}$ if there exists a family of vectors from a vector space over $\mathbb{F}$ whose linear independence relation is the same as the independence relation of the matroid. If the matroid is representable over any field, then it is called regular. A uniform matroid of rank $r$ over a ground set of size $n$ has rank function $r(X)=\min\{|X|,r\}$ and is denoted by $U_{r,n}$. Given pairwise disjoint sets $S_{1}\cup\dots\cup S_{q}\subseteq S$, the corresponding partition matroid $M=(S,r)$ is defined by setting $r(X)=|\{i\mid X\cap S_{i}\neq\emptyset\}|$ for all $X\subseteq S$ [^2] In the literature, partition matroids are often defined in a more general form where $r(X)=\sum_{i=1}^{q}\min\{|X\cap S_{i}|,g_{i}\}$ for some $g_{1},\dots,g_{q}\in\mathbb{Z}_{+}$. However, in this paper, we are interested only in the case when $g_{i}=1$ for all $i\in[q]$ except for at most one which is 00.. It is not difficult to see that both uniform and partition matroids are representable over the reals.
Submodular and polymatroid functions.
By $S$, we always denote a finite ground set. A set function $\varphi\colon 2^{S}\to\mathbb{R}$ is normalized if $\varphi(S)=1$. We say that $\varphi$ is increasing if $X\subseteq Y$ implies $\varphi(X)\leq\varphi(Y)$, and decreasing if $X\subseteq Y$ implies $\varphi(X)\geq\varphi(Y)$. The function $\varphi$ is submodular if
$\varphi(X)+\varphi(Y)\geq\varphi(X\cap Y)+\varphi(X\cup Y)$
for all $X,Y\subseteq S$, supermodular if $-\varphi$ is submodular, and modular if it is both sub- and supermodular.
A function $\varphi\colon 2^{S}\to\mathbb{R}_{+}$ is called a polymatroid function if it is an increasing, submodular function with $\varphi(\emptyset)=0$. Furthermore, if $\varphi(X)\leq k\cdot|X|$ holds for some $k\in\mathbb{R}_{+}$ and for every subset $X\subseteq S$, then $\varphi$ is a $k$-polymatroid function. Observe that polymatroid functions retain three of the four basic properties of matroid rank functions: (R1), (R2), and (R4), omitting only the subcardinality and integrality requirement. Thus, matroid rank functions form a subclass of polymatroid functions. As a sort of reverse statement, Helgason [21] showed that every integer-valued polymatroid function can be obtained as the quotient of a matroid rank function.
Proposition 2.1 (Helgason).
If $\varphi\colon 2^{S}\to\mathbb{Z}_{+}$ is an integer-valued polymatroid function over a ground set $S$, then there exists a matroid $M=(S^{\prime},r)$ and a mapping $\theta\colon S^{\prime}\to S$ such that $\varphi(X)=r(\theta^{-1}(X))$ for all $X\subseteq S$.
The base polyhedron of a nonnegative submodular function $\varphi\colon 2^{S}\to\mathbb{R}_{+}$ is defined as
$B(\varphi)\coloneqq\{x\in\mathbb{R}^{S}_{+}\mid x(S)=\varphi(S),~{}x(Z)\leq\varphi(Z)\text{ for every $Z\subseteq S$}\}.$
The following is a fundamental result in polyhedral combinatorics, see e.g. [18, Corollary 14.2.3].
Proposition 2.2.
If $\varphi$ is a nonnegative submodular function, then $B(\varphi)$ is non-empty. Furthermore, if $\varphi$ is integer-valued, then the vertices of $B(\varphi)$ are integer.
Coverage functions.
A coverage function $\varphi\colon 2^{S}\to\mathbb{R}_{+}$ is defined by a bipartite graph $G=(S,T;E)$ and weights $w\in\mathbb{R}^{T}_{+}$ by setting $\varphi(X)=\sum_{t\in N(X)}w_{t}$ for $X\subseteq S$, where $N(X)$ denotes the set of neighbours of $X$ in $G$. It is less known that coverage functions also have an alternative characterization involving inequalities [7]. Specifically, $\varphi$ is a coverage function if and only if $\varphi(\emptyset)=0$ and, for any $k\in\mathbb{Z}_{>0}$,
$\sum_{K\subseteq[k]}(-1)^{|K|}\varphi\bigl{(}A_{0}\cup\bigcup_{i\in K}A_{i}\bigr{)}\leq 0$ ($k$-Alt)
holds for any choice of subsets $A_{0},A_{1},\ldots,A_{k}\subseteq S$. It is worth noting that for $k=1$ and $k=2$, inequality ($k$-Alt) is equivalent with the property that $\varphi$ is increasing and increasing submodular, respectively. In general, for $k\geq 3$, the inequality describes a strengthened form of increasing submodularity [^3] In [7], it was shown that for functions over a finite domain, the class of coverage functions coincide with that of infinite-alternating functions, an important subclass of submodular functions that was studied by Choquet [10] in the analytic setting.. For any non-empty $A\subseteq S$, let $\varphi_{A}\colon 2^{S}\to\mathbb{R}$ be defined as
$\varphi_{A}(X)=\begin{cases}0&\text{if $X\cap A=\emptyset$,}\\ 1&\text{if $X\cap A\neq\emptyset$.}\end{cases}$ (1)
It is not difficult to see that $\varphi_{A}$ is a normalized coverage function. Choquet [10, Section 43] showed that these set functions correspond to the extreme rays of the convex cone of coverage functions; see also [30] for a different proof.
Proposition 2.3 (Choquet).
The set of normalized extremal elements of the convex cone of coverage functions over a finite set $S$ is $\{\varphi_{A}\mid\emptyset\neq A\subseteq S\}$.
For further details on coverage functions, we recommend [10, 30, 7].
From finite to infinite domain.
The above definitions extend naturally to set functions over an infinite domain. For clarity, we denote the ground set by $J$ rather than $S$ whenever it is infinite. In this context, let $(J,\mathcal{B})$ be a set algebra, i.e., $\mathcal{B}$ is a family of subsets of $J$ with $\emptyset\in\mathcal{B}$ and closed under taking complements and finite unions. Note that these imply that $\mathcal{B}$ is also closed under taking finite intersections. A set function $\varphi$ assigns a real value to each member of $\mathcal{B}$. By a slight abuse of notation, we call the members of $\mathcal{B}$ measurable. For set algebras $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$, a function $f\colon J_{1}\to J_{2}$ is measurable if $f^{-1}(B)\in\mathcal{B}_{1}$ for all $B\in\mathcal{B}_{2}$. For a set algebra $(J,\mathcal{B})$, we call a nonempty set $B\in\mathcal{B}$ an atom of the algebra if $A\subseteq B$, $A\in\mathcal{B}$ implies $A=\emptyset$ or $A=B$. We refer to a finitely additive measure, which is not necessarily nonnegative, as a charge. Properties of set functions, such as being normalized, increasing, decreasing, submodular, supermodular, modular and being a polymatroid can be generalized to functions over an infinite domain in a straightforward manner. Moreover, using the inequality-based definition, this is also doable for coverage functions. We call a set function finite if its domain is finite, and a set function is bounded, if the range of the set function is bounded.
Let $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$ be measurable spaces, and let $J=J_{1}\times J_{2}$. The product algebra $\mathcal{B}$ on $J$ is defined as the algebra generated by $\{X_{1}\times J_{2}\mid X_{1}\in\mathcal{B}_{1}\}\cup\{J_{1}\times X_{2}\mid X_{2}\in\mathcal{B}_{2}\}$, and it is denoted by $\mathcal{B}=\mathcal{B}_{1}\otimes\mathcal{B}_{2}$, see e.g. [17]. We refer to the measurable space $(J,\mathcal{B})$ thus obtained as the product of measurable spaces $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$. Let $\varphi_{1}\colon\mathcal{B}_{1}\to\mathbb{R}$ and $\varphi_{2}\colon\mathcal{B}_{2}\to\mathbb{R}$ be set functions. Then, $\varphi\colon\mathcal{B}\to\mathbb{R}$ is a coupling of $\varphi_{1}$ and $\varphi_{2}$ if $\varphi(X_{1}\times J_{2})=\varphi_{1}(X_{1})\cdot\varphi_{2}(J_{2})$ for every $X_{1}\in\mathcal{B}_{1}$ and $\varphi(J_{1}\times X_{2})=\varphi_{1}(J_{1})\cdot\varphi_{2}(X_{2})$ for every $X_{2}\in\mathcal{B}_{2}$. Similarly to the finite case, for $x\in J_{1}$, $y\in J_{2}$ and $Z\in\mathcal{B}$, we use $Z_{x}$ and $Z^{y}$ for denoting the $x$- and $y$-fibers of $Z$, and $\pi_{1},\pi_{2}$ denote the coordinate maps. By $\mathbbm{1}_{Z}$ we denote the indicator function of a set $Z$.
3 Coupling submodular functions
In this section we show that, unlike the tensor product, two matroids always admit a coupling. We first prove this statement for nonnegative submodular functions in Section 3.1 and then extend the discussion to polymatroid functions in Section 3.2. Matroids and the relation between coupling and tensor products are addressed in Section 3.3. We study $k$-alternating and coverage functions in Section 3.4. Finally, in Section 3.5 we explain how all these results can be extended to functions over an infinite ground set.
3.1 Submodular functions
As a first step, we verify that couplings exist for submodular functions. Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}_{+}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}_{+}$ be submodular functions over ground sets $S_{1}$ and $S_{2}$, and let $S=S_{1}\times S_{2}$. Recall that for any $Z\subseteq S$, $x\in S_{1}$ and $y\in S_{2}$, we use $Z_{x}=Z\cap(\{x\}\times S_{2})$ and $Z^{y}=Z\cap(S_{1}\times\{y\})$ to denote the $x$- and $y$-fibers of $Z$, respectively. For any $\mu_{1}\colon S_{1}\to\mathbb{R}_{+}$ and $\mu_{2}\colon S_{2}\to\mathbb{R}_{+}$, define the set function $b$ on $S$ by
$b(Z)\coloneqq\sum_{e_{1}\in S_{1}}\mu_{1}(e_{1})\cdot\varphi_{2}(\pi_{2}(Z_{e_{1}}))+\sum_{e_{2}\in S_{2}}\mu_{2}(e_{2})\cdot\varphi_{1}(\pi_{1}(Z^{e_{2}}))-\sum_{(e_{1},e_{2})\in Z}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2}).$ (2)
Note that, by the nonnegativity of $\varphi_{1}$ and $\varphi_{2}$, $\mu_{1}$ and $\mu_{2}$ can always be chosen such that $\mu_{1}(S_{1})=\varphi_{1}(S_{1})$ and $\mu_{2}(S_{2})=\varphi_{2}(S_{2})$. Therefore, the next theorem settles the existence of a coupling of two nonnegative submodular functions whenever $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$.
Theorem 3.1.
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}_{+}$ and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}_{+}$ be submodular functions over ground sets $S_{1}$ and $S_{2}$, respectively, satisfying $\varphi_{1}(\emptyset)=0$ and $\varphi_{2}(\emptyset)=0$. Furthermore, let $\mu_{1}\colon S_{1}\to\mathbb{R}_{+}$ and $\mu_{2}\colon S_{2}\to\mathbb{R}_{+}$ be such that $\mu_{1}(S_{1})=\varphi_{1}(S_{1})$ and $\mu_{2}(S_{2})=\varphi_{2}(S_{2})$. Then, the function $b$ defined in (2) is a submodular coupling of $\varphi_{1}$ and $\varphi_{2}$.
Proof.
Take arbitrary $X,Y\subseteq S_{1}\times S_{2}$, $e_{1}\in S_{1}$ and $e_{2}\in S_{2}$. Then we have
$\pi_{2}((X\cap Y)_{e_{1}})$ $=\pi_{2}(X_{e_{1}})\cap\pi_{2}(Y_{e_{1}}),$
$\pi_{2}((X\cup Y)_{e_{1}})$ $=\pi_{2}(X_{e_{1}})\cup\pi_{2}(Y_{e_{1}}),$
$\pi_{1}((X\cap Y)^{e_{2}})$ $=\pi_{1}(X^{e_{2}})\cap\pi_{1}(Y^{e_{2}}),$
$\pi_{1}((X\cup Y)^{e_{2}})$ $=\pi_{1}(X^{e_{2}})\cup\pi_{1}(Y^{e_{2}}).$
Together with the submodularity of $\varphi_{1}$ and $\varphi_{2}$ and the nonnegativity of $\mu_{1}$ and $\mu_{2}$, this implies that $b$ is a nonnegative combination of submodular functions minus a modular function, and is therefore submodular.
It remains to show that $b$ is a coupling. For each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$, we get
$b(Y_{1}\times Y_{2})$ $=\mu_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\mu_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})-\mu_{1}(Y_{1})\cdot\mu_{2}(Y_{2})$ (3)
$=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})-(\varphi_{1}(Y_{1})-\mu_{1}(Y_{1}))\cdot(\varphi_{2}(Y_{2})-\mu_{2}(Y_{2})).$ (4)
Thus, by $\mu_{1}(S_{1})=\varphi_{1}(S_{1})$ and $\mu_{2}(S_{2})=\varphi_{2}(S_{2})$, we have $b(Y_{1}\times Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$ whenever $Y_{1}=S_{1}$ or $Y_{2}=S_{2}$, concluding the proof of the theorem. ∎
It is worth emphasizing that if $\varphi_{1}$ and $\varphi_{2}$ are integer-valued, they admit an integer-valued submodular coupling. This follows by choosing $\mu_{1}$ and $\mu_{2}$ to be integer-valued in the proof.
3.2 Polymatroid functions
A noteworthy feature of the proof of Theorem 3.1 is that it relies on two arbitrarily chosen nonnegative modular functions satisfying $\mu_{1}(S_{1})=\varphi_{1}(S_{1})$ and $\mu_{2}(S_{2})=\varphi_{2}(S_{2})$. However, the resulting function $b$ may not be increasing, even if $\varphi_{1}$ and $\varphi_{2}$ are so. Therefore, to show that polymatroid functions admit a polymatroid coupling, $\mu_{1}$ and $\mu_{2}$ must be chosen carefully.
Theorem 3.2.
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}_{+}$ be a $k_{1}$-polymatroid and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}_{+}$ be a $k_{2}$-polymatroid function over ground sets $S_{1}$ and $S_{2}$, respectively. Then, $\varphi_{1}$ and $\varphi_{2}$ have a $(k_{1}\cdot k_{2})$-polymatroid coupling which is integer-valued if $\varphi_{1}$ and $\varphi_{2}$ are integer-valued.
Proof.
Let $\mu_{1}\in B(\varphi_{1})$ and $\mu_{2}\in B(\varphi_{2})$ be arbitrary elements of the base polyhedrons of $\varphi_{1}$ and $\varphi_{2}$, respectively, which are integers if $\varphi_{1}$ and $\varphi_{2}$ are integers. Note that such elements exist by Proposition 2.2. Let $b$ denote the submodular coupling of $\varphi_{1}$ and $\varphi_{2}$ defined as in (2) using $\mu_{1}$ and $\mu_{2}$.
Claim 3.3.
We have $b(\emptyset)=0$ and $b(Y_{1}\times Y_{2})\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$ for each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$, where equality holds if $\varphi_{1}(Y_{1})=\mu_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\mu_{2}(Y_{2})$.
Proof.
The equality $b(\emptyset)=0$ follows from $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$. For each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$, using (4), $\mu_{1}(Y_{1})\leq\varphi_{1}(Y_{1})$ and $\mu_{2}(Y_{2})\leq\varphi_{2}(Y_{2})$, we get
$b(Y_{1}\times Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})-(\varphi_{1}(Y_{1})-\mu_{1}(Y_{1}))\cdot(\varphi_{2}(Y_{2})-\mu_{2}(Y_{2}))\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2}).$
Here equality holds if $\varphi_{1}(Y_{1})=\mu_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\mu_{2}(Y_{2})$. ∎
The function $b$ thus obtained is still not necessarily increasing. To get an increasing set function, we define $\varphi$ on $S$ by setting
$\varphi(Z)\coloneqq\min\{b(Z^{\prime})\mid Z^{\prime}\supseteq Z\}.$
The following technical claim is folklore, see e.g. [18]. We include a proof to make the paper self-contained.
Claim 3.4.
$\varphi$ is an increasing submodular function over $S_{1}\times S_{2}$.
Proof.
Clearly, $\varphi$ is increasing by definition. To prove submodularity, let $Z_{1},Z_{2}\subseteq S_{1}\times S_{2}$. Then, for $i=1,2$, there exists $Z^{\prime}_{i}\supseteq Z_{i}$ such that $\varphi(Z_{i})=b(Z^{\prime}_{i})$. Then, using the submodularity of $b$ and the definition of $\varphi$, we get
$\varphi(Z_{1})+\varphi(Z_{2})$ $=b(Z^{\prime}_{1})+b(Z^{\prime}_{2})$
$\geq b(Z^{\prime}_{1}\cap Z^{\prime}_{2})+b(Z^{\prime}_{1}\cup Z^{\prime}_{2})$
$\geq\varphi(Z_{1}\cap Z_{2})+\varphi(Z_{1}\cup Z_{2}),$
proving the claim. ∎
Finally, we show that $\varphi$ and $b$ coincide on product sets.
Claim 3.5.
$\varphi(Y_{1}\times Y_{2})=b(Y_{1}\times Y_{2})$ holds for each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$.
Proof.
We need to show that $b(Z)\geq b(Y_{1}\times Y_{2})$ holds if $Y_{1}\times Y_{2}\subseteq Z\subseteq S_{1}\times S_{2}$. Using that $\varphi_{2}$ is increasing and $\mu_{2}(X)\leq\varphi_{2}(X)$ holds for any $X\subseteq S_{2}$, we get
$\sum_{e_{1}\in S_{1}}\mu_{1}(e_{1})\cdot\varphi_{2}(\pi_{2}(Z_{e_{1}}))$ $\geq\sum_{e_{1}\in Y_{1}}\mu_{1}(e_{1})\cdot\varphi_{2}(Y_{2})+\sum_{e_{1}\in S_{1}\setminus Y_{1}}\mu_{1}(e_{1})\cdot\varphi_{2}(\pi_{2}(Z_{e_{1}}))$
$\geq\sum_{e_{1}\in Y_{1}}\mu_{1}(e_{1})\cdot\varphi_{2}(Y_{2})+\sum_{e_{1}\in S_{1}\setminus Y_{1}}\mu_{1}(e_{1})\cdot\mu_{2}(\pi_{2}(Z_{e_{1}}))$
$=\mu_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\sum_{\begin{subarray}{c}(e_{1},e_{2})\in Z\setminus(Y_{1}\times S_{2})\end{subarray}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2}).$ (5)
Similarly, using that $\varphi_{1}$ is increasing and $\mu_{1}(X)\leq\varphi_{1}(X)$ holds for any $X\subseteq S_{1}$, we get
$\sum_{e_{2}\in S_{2}}\mu_{2}(e_{2})\cdot\varphi_{1}(\pi_{1}(Z^{e_{2}}))$ $\geq\sum_{e_{2}\in Y_{2}}\mu_{2}(e_{2})\cdot\varphi_{1}(Y_{1})+\sum_{e_{2}\in S_{2}\setminus Y_{2}}\mu_{2}(e_{2})\cdot\varphi_{1}(\pi_{1}(Z^{e_{2}}))$
$\geq\sum_{e_{2}\in Y_{2}}\mu_{2}(e_{2})\cdot\varphi_{1}(Y_{1})+\sum_{e_{2}\in S_{2}\setminus Y_{2}}\mu_{2}(e_{2})\cdot\mu_{1}(\pi_{1}(Z^{e_{2}}))$
$=\mu_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})+\sum_{\begin{subarray}{c}(e_{1},e_{2})\in Z\setminus(S_{1}\times Y_{2})\end{subarray}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2}).$ (6)
Using the definition (2) of $b$, the equality (3), and the sum of the inequalities (5) and (6), we get
$b(Z)$ $\geq\mu_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\sum_{\begin{subarray}{c}(e_{1},e_{2})\in Z\setminus(Y_{1}\times S_{2})\end{subarray}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2})+\mu_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})+\sum_{\begin{subarray}{c}(e_{1},e_{2})\in Z\setminus(S_{1}\times Y_{2})\end{subarray}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2})$
$~{}~{}~{}~{}-\sum_{(e_{1},e_{2})\in Z}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2})$
$=\mu_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\mu_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})+\sum_{\begin{subarray}{c}(e_{1},e_{2})\in Z\\ e_{1}\not\in Y_{1},e_{2}\not\in Y_{2}\end{subarray}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2})-\sum_{(e_{1},e_{2})\in Y_{1}\times Y_{2}}\mu_{1}(e_{1})\cdot\mu_{2}(e_{2})$
$\geq\mu_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\mu_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})-\mu_{1}(Y_{1})\cdot\mu_{2}(Y_{2})$
$=b(Y_{1}\times Y_{2}).$
This concludes the proof of the claim. ∎
By Claim 3.4, $\varphi$ is increasing submodular. Claim 3.5 implies that $\varphi(\emptyset)=b(\emptyset)=0$, thus $\varphi\geq 0$. Using $\varphi_{1}(S_{1})=\mu_{1}(S_{1})$, $\varphi_{2}(S_{2})=\mu_{2}(S_{2})$, and Claims 3.3 and 3.5, we get that $\varphi(Y_{1}\times S_{2})=b(Y_{1}\times S_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(S_{2})$ and $\varphi(S_{1}\times Y_{2})=b(S_{1}\times Y_{2})=\varphi_{1}(S_{1})\times\varphi_{2}(Y_{2})$ holds for $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$, that is, $\varphi$ is a coupling of $\varphi_{1}$ and $\varphi_{2}$. For $(e_{1},e_{2})\in S_{1}\times S_{2}$, by Claims 3.3 and 3.5, we have
$\varphi(\{(e_{1},e_{2})\})=b(\{(e_{1},e_{2})\})\leq\varphi_{1}(\{e_{1}\})\cdot\varphi_{2}(\{e_{2}\})\leq k_{1}\cdot k_{2},$
that is, $\varphi$ is a $(k_{1}\cdot k_{2})$-polymatroid function. Finally, if $\varphi_{1}$ and $\varphi_{2}$ are integer-valued, then $\varphi$ is integer-valued as well by construction. This finishes the proof of the theorem. ∎
3.3 Matroids and tensor products
One of the main motivations for our work was to extend the notion of coupling to matroids. For ease of discussion, we say that $M=(S_{1}\times S_{2},r)$ is a coupling of matroids $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$ if $r$ is a coupling of $r_{1}$ and $r_{2}$. In Section 3.3.1, we show that matroids always admit a coupling that nearly satisfies the requirements of being a tensor product. Our construction, as it turns out, has deep connections to amalgams of matroids. Then, in Section 3.3.2, we characterize couplings that are tensor products as well. Based on our observations, in Section 3.3.3, we highlight a surprising phenomenon: the linearity of a matroid is closely related to whether it admits a tensor product with the uniform matroid $U_{2,3}$. This result could serve as a useful tool in studying linear matroids.
3.3.1 Matroid couplings and amalgams
Since the rank function of any matroid is an integer-valued $1$-polymatroid function, Theorem 3.2 can be applied. Actually, the proof implies an even stronger result.
Corollary 3.6.
Let $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$ be matroids, let $B_{1}$ be a basis of $M_{1}$ and $B_{2}$ be a basis of $M_{2}$, and set $S\coloneqq S_{1}\times S_{2}$. Then, $M_{1}$ and $M_{2}$ have a coupling $M=(S,r)$ such that $r(Y_{1}\times Y_{2})=r_{1}(Y_{1})\cdot r_{2}(Y_{2})$ whenever $Y_{1}\subseteq B_{1}$ or $Y_{2}\subseteq B_{2}$.
Proof.
Let $\mu_{1}\coloneqq\chi_{B_{1}}$ and $\mu_{2}\coloneqq\chi_{B_{2}}$ be the characteristic vectors of $B_{1}$ and $B_{2}$, respectively. Clearly, $\chi_{i}\in B(r_{i})$ holds for $i=1,2$. By Theorem 3.2, $M_{1}$ and $M_{2}$ admit a coupling $M=(S,r)$, where
$r(Z)=\min_{W\supseteq Z}\left\{\sum_{e_{1}\in B_{1}}r_{2}(\pi_{2}(W_{e_{1}}))+\sum_{e_{2}\in B_{2}}r_{1}(\pi_{1}(W^{e_{2}}))-|W\cap(B_{1}\times B_{2})|\right\}.$
Let $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$. By Claim 3.5, the minimum is attained on $W=Z$ for $Z=Y_{1}\times Y_{2}$. Furthermore, if $Y_{1}\subseteq B_{1}$ or $Y_{2}\subseteq B_{2}$, then $\varphi_{1}(Y_{1})=\mu_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\mu_{2}(Y_{2})$, and thus $r(Y_{1}\times Y_{2})=r_{1}(Y_{1})\cdot r_{2}(Y_{2})$ by Claim 3.3. This completes the proof of the corollary. ∎
Corollary 3.6 has a somewhat surprising implication: for a fixed rank $r$ and ground set size $n$, there exists a function defined on a finite ground set that contains every rank-$r$ matroid in some sense. For $r,n\in\mathbb{Z}_{+}$, let $N_{r,n}$ denote the number of rank-$r$ matroids on $n$ elements up to isomorphism.
Corollary 3.7.
There exists a matroid $M$ over $[n]^{N_{r,n}}$ of rank $r^{N_{r,n}}$ such that any rank-$r$ matroid on $[n]$ is a projection of $M$ to one of its coordinates.
Proof.
The results follows by applying the coupling of matroid pairs iteratively for every rank-$r$ matroid over a ground set of size $n$. ∎
Recall that two matroids do not necessarily admit a tensor product. However, the main message of Corollary 3.6 is that there always exists a coupling that locally satisfies the properties of the tensor product in the sense that $r(Y_{1}\times Y_{2})=r_{1}(Y_{1})\cdot r_{2}(Y_{2})$ holds, not universally for all $Y_{1},Y_{2}$, but specifically when at least one of $Y_{1}\subseteq B_{1}$ or $Y_{2}\subseteq B_{2}$ is satisfied. In particular, this implies that $M|(S_{1}\times\{y\})$ is isomorphic to $M_{1}$ for each $y\in B_{2}$ by the natural bijection to $S_{1}$, and $M|(\{x\}\times S_{2})$ is isomorphic to $M_{2}$ for each $x\in B_{1}$ by the natural bijection to $S_{2}$.
Remark 3.8.
In the previous discussion, our strategy was to verify the existence of a coupling first for polymatroid functions and then for matroids. However, one could also take a reverse approach. The construction in Theorem 3.2 can be formalized directly for matroids, as demonstrated in the proof of Corollary 3.6. Since every integer-valued polymatroid function is the quotient of a matroid rank function by Proposition 2.1, this immediately implies the existence of a coupling for integer-valued polymatroid functions. Extending this, one can readily establish the statement for rational-valued polymatroid functions. Finally, by approximating a real-valued polymatroid function with rational-valued ones, the existence of a coupling for general polymatroid functions follows.
Let $N_{0}=(T_{1}\cap T_{2},r_{0}^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{1})$ and $N_{2}=(T_{2},r^{\prime}_{2})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. A matroid $N=(T,r)$ is called an amalgam of $N_{1}$ and $N_{2}$ if $N|T_{1}=N_{1}$ and $N|T_{2}=N_{2}$. An amalgam $N$ is called free if for any amalgam $N^{\prime}$, each independent set of $N^{\prime}$ is independent in $N$. The following result appears, for example, in [33, Proposition 11.4.2].
Proposition 3.9.
Let $N_{0}=(T_{1}\cap T_{2},r^{\prime})$, $N_{1}=(T_{1},r^{\prime}_{1})$, and $N_{2}=(T_{2},r^{\prime}_{2})$ be matroids such that $N_{1}|(T_{1}\cap T_{2})=N_{2}|(T_{1}\cap T_{2})=N_{0}$, and let $T\coloneqq T_{1}\cup T_{2}$. Define $r\colon 2^{T}\to\mathbb{Z}_{+}$ by
$r(Z)\coloneqq\min\{r^{\prime}_{1}(Y\cap T_{1})+r^{\prime}_{2}(Y\cap T_{2})-r^{\prime}(Y\cap T_{1}\cap T_{2})\mid Y\supseteq Z\}.$
If $N_{0}$ is a modular matroid, then $r$ is submodular. Moreover, if $r$ is submodular, then $N=(T,r)$ is a free amalgam of $N_{1}$ and $N_{2}$.
If the function $r$ defined in Proposition 3.9 is submodular, then $N=(T,r)$ is called the proper amalgam of $N_{1}$ and $N_{2}$. Both free and proper amalgams are unique, if they exist. It turns out, though, that for a given pair of matroids, there may be no free amalgam, or the free amalgam may not be proper, or no amalgams may exist at all.
Given matroids $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$, a basis $B_{1}$ of $M_{1}$, and a basis $B_{2}$ of $M_{2}$, consider the matroids $N_{1}=\bigoplus_{e_{2}\in B_{2}}M_{1}$ on ground set $S_{1}\times B_{2}$ and $N_{2}=\bigoplus_{e_{1}\in B_{1}}M_{2}$ on ground set $B_{1}\times S_{2}$. Observe that $(S_{1}\times B_{2})\cap(B_{1}\times S_{2})=B_{1}\times B_{2}$ and $N_{1}|(B_{1}\times B_{2})=N_{2}|(B_{1}\times B_{2})$ is the free matroid on $B_{1}\times B_{2}$. Therefore, Proposition 3.9 implies that the proper amalgam of $N_{1}$ and $N_{2}$ exists. Our construction for the coupling of $M_{1}$ and $M_{2}$ shown in the proof of 3.6 can be obtained from the proper amalgam of $N_{1}$ and $N_{2}$ by adding the elements of $(S_{1}\setminus B_{1})\times(S_{2}\setminus B_{2})$ as loops. We prove that the same construction gives a coupling for any amalgam of $N_{1}$ and $N_{2}$.
Theorem 3.10.
Let $M_{1}=(S_{1},r_{1})$ and $M_{2}=(S_{2},r_{2})$ be matroids, and let $B_{1}$ be a basis of $M_{1}$ and $B_{2}$ be a basis of $M_{2}$. Consider the matroids $N_{1}=\bigoplus_{e_{2}\in B_{2}}M_{1}$ on ground set $S_{1}\times B_{2}$ and $N_{2}=\bigoplus_{e_{1}\in B_{1}}M_{2}$ on ground set $B_{1}\times S_{2}$, and let $S\coloneqq S_{1}\times S_{2}$. If $M=(S,r)$ is a matroid such that each each element of $(S_{1}\setminus B_{1})\times(S_{2}\setminus B_{2})$ is a loop in $M$ and $M|((S_{1}\times B_{2})\cup(B_{1}\times S_{2}))$ is an amalgam of $N_{1}$ and $N_{2}$, then $M$ is a coupling of $M_{1}$ and $M_{2}$.
Proof.
We show that $r(Y_{1}\times S_{2})=r_{1}(Y_{1})\cdot r_{2}(S_{2})$ holds for each $Y_{1}\subseteq S_{1}$; a similar reasoning proves that $r(S_{1}\times Y_{2})=r_{1}(S_{1})\cdot r_{2}(Y_{2})$ holds for each $Y_{2}\subseteq S_{2}$. Since $B_{1}\cap Y_{1}$ is independent in $M_{1}$, it can be extended to a maximum independent set $I_{1}$ of $Y_{1}$ with $|I_{1}|=r_{1}(Y_{1})$. Since $I_{1}\times B_{2}$ is independent in $N_{1}$, it is independent in $N$, thus $r(Y_{1}\times S_{2})\geq|I_{1}\times B_{2}|=r_{1}(Y_{1})\cdot r_{2}(S_{2})$. To see the reverse inequality, observe that $I_{1}\times B_{2}$ spans $Y_{1}\times B_{2}$ in $N_{1}$, $(B_{1}\cap Y_{1})\times B_{2}\subseteq I_{1}\times B_{2}$ spans $(B_{1}\cap Y_{1})\times S_{2}$ in $N_{2}$, and each element of $(Y_{1}\setminus B_{1})\times(S_{2}\setminus B_{2})$ is a loop of $N$. Thus $I_{1}\times B_{2}$ spans $Y_{1}\times S_{2}$ in $N$, implying $r(Y_{1}\times S_{2})\leq|I_{1}\times B_{2}|=r_{1}(Y_{1})\cdot r_{2}(S_{2})$. ∎
3.3.2 Tensor products
From the definitions alone, it is clear that coupling and the tensor product are closely related concepts. The following theorem characterizes which couplings satisfy the stronger properties of the tensor product.
Theorem 3.11.
Let $M_{1}=(S_{1},r_{1})$, $M_{2}=(S_{2},r_{2})$ and $M=(S,r)$ be matroids such that $S=S_{1}\times S_{2}$. Then, the following are equivalent:
  1. (i)
    $r(Y_{1}\times Y_{2})=r_{1}(Y_{1})\cdot r_{2}(Y_{2})$ for each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$,
  2. (ii)
    $r(\{e_{1}\}\times Y_{2}))=r_{1}(e_{1})\cdot r_{2}(Y_{2})$ for each $Y_{2}\subseteq S_{2}$ and $e_{1}\in S_{1}$, $r(Y_{1}\times\{e_{2}\}))=r_{1}(Y_{1})\cdot r_{2}(e_{2})$ for each $Y_{1}\subseteq S_{1}$ and $e_{2}\in S_{2}$, and $r(S)=r_{1}(S_{1})\cdot r_{2}(S_{2})$,
  3. (iii)
    $M$ is a coupling of $M_{1}$ and $M_{2}$ and $r((e_{1},e_{2}))=r_{1}(e_{1})\cdot r_{2}(e_{2})$ for each $(e_{1},e_{2})\in S$.
Proof.
The equivalence of (i) and (ii) was shown by Las Vergnas [26], and it is also clear that (i) implies (iii). Assume now that (iii) holds, we prove that it implies (ii). For the proof, we will use the following simple observation.
Claim 3.12.
If $J_{1}\subseteq S_{1}$ is independent in $M_{1}$ and $J_{2}\subseteq S_{2}$ is independent in $M_{2}$, then
$M|(J_{1}\times S_{2})=\bigoplus\limits_{e_{1}\in J_{1}}M|(\{e_{1}\}\times S_{2})\text{ and }M|(S_{1}\times J_{2})=\bigoplus\limits_{e_{2}\in J_{2}}M|(S_{1}\times\{e_{2}\}).$
Proof.
Using that $M$ is a coupling, we get that
$\sum_{e_{1}\in J_{1}}r(\{e_{1}\}\times S_{2})=\sum_{e_{1}\in J_{1}}r_{1}(e_{1})\cdot r_{2}(S_{2})=|J_{1}|\cdot r_{2}(S_{2})=r_{1}(J_{1})\cdot r_{2}(S_{2})=r(J_{1}\times S_{2}),$
implying $M|(J_{1}\times S_{2})=\bigoplus_{e_{1}\in J_{1}}M|(\{e_{1}\}\times S_{2})$. The second statement follows similarly. ∎
First we show that $r(Y_{1}\times Y_{2})\geq r_{1}(Y_{1})\cdot r_{2}(Y_{2})$ holds for each $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$. Let $I_{1}\subseteq Y_{1}$ and $I_{2}\subseteq Y_{2}$ be maximal independent sets of $M_{1}$ and $M_{2}$ contained in $Y_{1}$ and $Y_{2}$, respectively. 3.12 implies that each circuit contained in $I_{1}\times S_{2}$ intersects at most one of the sets of the form $\{e_{1}\}\times S_{2}$ for $e_{1}\in I_{1}$. Similarly, each circuit contained in $S_{1}\times I_{2}$ intersects at most one of the sets of the form $S_{1}\times\{e_{2}\}$ for $e_{2}\in I_{2}$. These show that each circuit in $I_{1}\times I_{2}$ is a loop. Since $r((e_{1},e_{2}))=r_{1}(e_{1})\cdot r_{2}(e_{2})=1$ holds for $(e_{1},e_{2})\in I_{1}\times I_{2}$ by assumption, the set $I_{1}\times I_{2}$ does not contain any loops. This proves that $I_{1}\times I_{2}$ is independent in $M$, thus $r(Y_{1}\times Y_{2})\geq r_{1}(Y_{1})\cdot r_{2}(Y_{2})$.
Next we show that $r(\{e_{1}\}\times Y_{2})=r_{1}(e_{1})\cdot r_{2}(Y_{2})$ holds for each $e_{1}\in S_{1}$ and $Y_{2}\subseteq S_{2}$. If $e_{1}$ is a loop in $M_{1}$, then $r(\{e_{1}\}\times S_{2})=r_{1}(e_{1})\cdot r_{2}(S_{2})=0$, thus $r(\{e_{1}\}\times Y_{2})=0$. Otherwise, let $B_{1}$ be a basis of $M_{1}$ containing $e_{1}$. By the previous paragraph, we have $r(\{b_{1}\}\times Y_{2})\geq r_{2}(Y_{2})$ for $b_{1}\in B_{1}$. Furthermore, $M|(B_{1}\times Y_{2})=\bigoplus_{b_{1}\in B_{1}}M|(\{b_{1}\}\times Y_{2})$ by 3.12. Thus,
$r(S_{1}\times Y_{2})\geq r(B_{1}\times Y_{2})=\sum_{b_{1}\in B_{1}}r(\{b_{1}\}\times Y_{2})\geq\sum_{b_{1}\in B_{1}}r_{2}(Y_{2})=r_{1}(S_{1})\cdot r_{2}(Y_{2})=r(S_{1}\times Y_{2}),$
where the last equality holds by the assumption that $M$ is a coupling. Therefore, equality holds throughout, implying that $r(\{b_{1}\}\times Y_{2})=r_{2}(Y_{2})$ for each $b_{1}\in B_{1}$. In particular, this shows that $r(\{e_{1}\}\times Y_{2})=r_{1}(e_{1})\cdot r_{2}(Y_{2})$. Similarly, we get that $r(Y_{1}\times\{e_{2}\})=r_{1}(Y_{1})\cdot r_{2}(e_{2})$ for each $Y_{1}\subseteq S_{1}$ and $e_{2}\in S_{2}$. Finally, $r(S)=r_{1}(S_{1})\cdot r_{2}(S_{2})$ follows from the assumption that $M$ is a coupling. ∎
If $e_{1}$ or $e_{2}$ is a loop in the corresponding matroid, then $(e_{1},e_{2})$ is a loop in any coupling by definition. Hence if $(e_{1},e_{2})$ is not a loop in a coupling, then $r((e_{1},e_{2}))=r_{1}(e_{1})\cdot r_{2}(e_{2})$ holds. Therefore, an interesting consequence of Theorem 3.11 is the following.
Corollary 3.13.
Any loopless coupling of two matroids is also a tensor product.
3.3.3 Linear matroids and Ingleton’s inequality
Tools for determining the linear representability of a matroid include the use of excluded minors and Ingleton’s inequality [22], who showed that the rank function of any linear matroid $M=(S,r)$ satisfies
$r(A\cup B)+r(A\cup C)+r(A\cup D)+r(B\cup C)+r(B\cup D)\geq$
$r(A)+r(B)+r(A\cup B\cup C)+r(A\cup B\cup D)+r(C\cup D)$ (7)
for every $A,B,C,D\subseteq S$. In this sense, the inequality describes a necessary condition for linearity. Since the uniform matroid $U_{2,3}$ is regular, it has a matroid tensor product with any linear matroid. In other words, for a matroid, the existence of a matroid tensor product with $U_{2,3}$ is also a necessary condition for linearity. Our next result shows that this property does not only imply Ingleton’s inequality for polymatroid functions but also yields a slightly weaker form for general submodular functions.
Theorem 3.14.
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ be a submodular function and let $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}_{+}$ denote the rank function of the uniform matroid $U_{2,3}$.
  1. (a)
    If $\varphi_{1}$ and $\varphi_{2}$ have a submodular tensor product, then $\varphi_{1}$ satisfies Ingleton’s inequality (7) for all pairwise disjoint sets $A,B,C,D\subseteq S_{1}$.
  2. (b)
    If $\varphi_{1}$ and $\varphi_{2}$ have a tensor product which is a polymatroid function, then $\varphi_{1}$ satisfies Ingleton’s inequality (7) for all sets $A,B,C,D\subseteq S_{1}$.
Proof.
Let $A,B,C,D\subseteq S_{1}$, $S\coloneqq S_{1}\times S_{2}$, and $\varphi\colon 2^{S}\to\mathbb{R}$ be a submodular tensor product of $\varphi_{1}$ and $\varphi_{2}$ which is increasing when proving (b). We denote the elements of $S_{2}$ by $1$, $2$, and $3$. To improve readability, we use several abbreviations in the following inequalities: for $H=\{h_{1},\dots,h_{k}\}\subseteq\{1,2,3\}$, we use $h_{1}\dots h_{k}$ to denote $H$, for $X\in\{A,B,C,D\}$, we let $X_{H}$ denote $X\times H$. Lastly, we omit the $\cup$ symbols, e.g., $AB\times 12$ denotes $(A\cup B)\times\{1,2\}$ and $A_{12}B_{13}$ denotes $(A\times\{1,2\})\cup(B\times\{1,3\})$. Consider the following inequalities.
$\varphi(AB\times 2)+\varphi(A_{123}C_{1})$ $\geq\varphi(A\times 2)+\varphi(A_{123}B_{2}C_{1})$
$\varphi(AC\times 1)+\varphi(A\times 123)$ $\geq\varphi(A\times 1)+\varphi(A_{123}C_{1})$
$\varphi(AD\times 3)+\varphi(A_{3}B_{2}C_{1})$ $\geq\varphi(A\times 3)+\varphi(A_{3}B_{2}C_{1}D_{3})$
$\varphi(BD\times 2)+\varphi(A_{3}B_{2}C_{1}D_{3})$ $\geq\varphi(B\times 2)+\varphi(A_{3}B_{2}C_{1}D_{23})$
$\varphi(ABC\times 3)+\varphi(BC\times 123)$ $\geq\varphi(BC\times 3)+\varphi(A_{3}B_{123}C_{123})$
$\varphi(ABCD\times 1)+\varphi(A_{3}B_{2}C_{1}D_{123})$ $\geq\varphi(CD\times 1)+\varphi(A_{13}B_{12}C_{1}D_{123})$
$\varphi(ABCD\times 2)+\varphi(A_{123}B_{12}C_{1}D_{123})$ $\geq\varphi(ABC\times 2)+\varphi(A_{123}B_{12}C_{12}D_{123})$
$\varphi(A\times 123)+\varphi(A_{13}B_{12}C_{1}D_{123})$ $\geq\varphi(A\times 13)+\varphi(A_{123}B_{12}C_{1}D_{123})$
$\varphi(A_{123}B_{2}C_{1})+\varphi(A_{3}B_{123}C_{123})$ $\geq\varphi(A_{3}B_{2}C_{1})+\varphi(ABC\times 123)$
$\varphi(D\times 123)+\varphi(A_{3}B_{2}C_{1}D_{23})$ $\geq\varphi(D\times 23)+\varphi(A_{3}B_{2}C_{1}D_{123})$
$\varphi(BC\times 123)+\varphi(A_{123}B_{12}C_{12}D_{123})$ $\geq\varphi(BC\times 12)+\varphi(ABCD\times 123)$
If $A$, $B$, $C$, and $D$ are pairwise disjoint, then each of the inequalities has form $\varphi(Y_{1})+\varphi(Y_{2})\geq\varphi(Y_{1}\cap Y_{2})+\varphi(Y_{1}\cup Y_{2})$ for some $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$, thus each of them holds by the submodularity of $\varphi$. In general, if they are not necessarily pairwise disjoint, then each of them has form $\varphi(Y_{1})+\varphi(Y_{2})\geq\varphi(Z)+\varphi(Y_{1}\cup Y_{2})$ for some $Y_{1}\subseteq S_{1}$, $Y_{2}\subseteq S_{2}$ and $Z\subseteq Y_{1}\cap Y_{2}$, thus each of them holds if $\varphi$ is a polymatroid function. Therefore, for proving statements (a) and (b), we may assume that each of these eleven inequalities holds.
In each of the inequalities, substitute $\varphi(Y_{1}\times Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})=\varphi_{1}(Y_{1})\cdot\min\{|Y_{2}|,2\}$ for each set of the form $Y_{1}\times Y_{2}$ for $Y_{1}\subseteq S_{1}$ and $Y_{2}\subseteq S_{2}$. Then, from the sum of above eleven inequalities, we obtain
$\varphi_{1}(AB)+\varphi_{1}(AC)+\varphi_{1}(AD)+\varphi_{1}(BC)+\varphi_{1}(BD)$ $\geq\varphi_{1}(A)+\varphi_{1}(B)+2\varphi_{1}(ABC)+\varphi_{1}(CD).$ (8)
Similarly, by the symmetry of the roles of $C$ and $D$, we obtain
$\varphi_{1}(AB)+\varphi_{1}(AC)+\varphi_{1}(AD)+\varphi_{1}(BC)+\varphi_{1}(BD)$ $\geq\varphi_{1}(A)+\varphi_{1}(B)+2\varphi_{1}(ABD)+\varphi_{1}(CD).$ (9)
The sum of (8) and (9) yields Ingleton’s inequality (7). ∎
It was shown by Las Vergnas [26] that the Vámos matroid and the uniform matroid $U_{2,3}$ do not have a matroid tensor product. The Vámos matroid is the matroid on ground set $\{a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},d_{1},d_{2}\}$ in which each subset of size 4 is a basis except for $A\cup B$, $A\cup C$, $A\cup D$, $B\cup C$ and $B\cup D$, where $A=\{a_{1},a_{2}\}$, $B=\{b_{1},b_{2}\}$, $C=\{c_{1},c_{2}\}$ and $D=\{d_{1},d_{2}\}$. As the rank function of the Vámos matroid does not satisfy Ingleton’s inequality (7) for the pairwise disjoint sets $A$, $B$, $C$, and $D$, 3.14 implies the following strengthening of the result of Las Vergnas [26].
Corollary 3.15.
The rank function of the Vámos matroid does not have a submodular tensor product with the rank function of the uniform matroid $U_{2,3}$.
For any fixed ground set $S_{1}$ and fixed matroid $N=(S_{2},r_{2})$, the set of submodular (or polymatroid) functions $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ which have a tensor product with $r_{2}$ forms a polyhedron in $\mathbb{R}^{2^{S_{1}}}$, that is, they can be described as functions $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}$ satisfying certain linear inequalities. If we choose $N$ to be a regular matroid, then each of these inequalities is a linear rank inequality, that is, a linear inequality satisfied by the rank function of any representable matroid. Ingleton’s inequality (7) was the first known linear rank inequality, and since then infinite families of independent linear rank inequalities has been found [12, 24]. Interestingly, all currently known linear rank inequalities follow from the so-called common information property of the rank function of representable matroids [12, 4], though other linear inequalities have been found that are satisfied by rank functions of matroids representable over fields of certain characteristics but not by all representable matroids [13]. For defining the common information property, let us call every polymatroid function 00-CI-compliant, and call a polymatroid function $f\colon 2^{S}\to\mathbb{R}$ is $k$-CI-compliant if for any subsets $A,B\subseteq S$, there exists a $(k-1)$-compliant function $f^{\prime}\colon 2^{S^{\prime}}\to\mathbb{R}$ and a subset $Z\subseteq S^{\prime}$ such that $S\subseteq S^{\prime}$, $f^{\prime}(X)=f(X)$ for all $X\subseteq S$, $f^{\prime}(A\cup Z)=f(A)$, $f^{\prime}(B\cup Z)=f(B)$, and $f^{\prime}(Z)=f(A)+f(B)-f(A\cup B)$. Then, a polymatroid satisfies the common information property if it is $k$-CI-compliant for each integer $k\geq 0$. Note that the common information property is not sufficient for the representability of a matroid, e.g. all rank-3 matroids satisfy it [4] while e.g. the non-Desargues matroid is not representable [22]. It would be interesting to see the relations between the class of polymatroid functions having tensor products with $U_{2,3}$, the class of polymatroid functions having tensor product with all regular matroids, and the class of polymatroid functions satisfying the common information property.
3.4 Coverage and $k$-alternating functions
Coverage functions naturally arise in applications like set cover, influence maximization, sensor placement, feature selection, and data summarization. The goal of this section is to show that if one of the functions is a coverage function, then there exist couplings with stronger properties. For some $k\in\mathbb{Z}_{>0}$, let us call a function $\varphi\colon 2^{S}\to\mathbb{R}_{+}$ $k$-alternating if $\varphi(\emptyset)=0$ and it satisfies ($k$-Alt) for the given $k$.
Theorem 3.16.
Let $\varphi_{1}\colon 2^{S_{1}}\to\mathbb{R}_{+}$ be a coverage function and $\varphi_{2}\colon 2^{S_{2}}\to\mathbb{R}_{+}$ be a $k$-alternating function over ground sets $S_{1}$ and $S_{2}$, respectively. Then, $\varphi_{1}$ and $\varphi_{2}$ have a $k$-alternating tensor product which is integer-valued if $\varphi_{1}$ and $\varphi_{2}$ are integer-valued.
Proof.
Assume first that $\varphi_{1}=\varphi_{A}$ for some $\emptyset\neq A\subseteq S_{1}$ as defined in (1). We define the tensor product of $\varphi_{A}$ and $\varphi_{2}$ to be
$(\varphi_{A}\otimes\varphi_{2})(X)\coloneqq\varphi_{2}(\pi_{2}((A\times S_{2})\cap X))$
for any $X\subseteq S_{1}\times S_{2}$. For each pair of sets $Y_{1}\subseteq S_{1}$, $Y_{2}\subseteq S_{2}$, we have $(\varphi_{A}\otimes\varphi_{2})(Y_{1}\times Y_{2})=\varphi_{2}(Y_{2})$ if $A\cap Y_{1}\neq\emptyset$ and $(\varphi_{A}\otimes\varphi_{2})(Y_{1}\times Y_{2})=0$ if $A\cap Y_{1}=\emptyset$, thus $\varphi_{A}\otimes\varphi_{2}$ is indeed a tensor product. We claim that it is $k$-alternating as well. To see this, let $X_{0},X_{1},\dots,X_{k}\subseteq S_{1}\times S_{2}$. Then, the $k$-alternating condition ($k$-Alt) requires that
$\sum_{K\subseteq[k]}(-1)^{|K|}(\varphi_{A}\otimes\varphi_{2})\bigl{(}X_{0}\cup\bigcup_{i\in K}X_{i}\bigr{)}$ $=\sum_{K\subseteq[k]}(-1)^{|K|}\varphi_{2}\Bigl{(}\pi_{2}{\bigl{(}}(A\times S_{2})\cap(X_{0}\cup\bigcup_{i\in K}X_{i})\bigr{)}\Bigr{)}$
$=\sum_{K\subseteq[k]}(-1)^{|K|}\varphi_{2}\Bigl{(}\pi_{2}{\bigl{(}}(A\times S_{2})\cap X_{0}\bigr{)}\cup\bigcup_{i\in K}\pi_{2}{\bigl{(}}(A\times S_{2})\cap X_{i})\bigr{)}\Bigr{)}$
$=\sum_{K\subseteq[k]}(-1)^{|K|}\varphi_{2}\bigl{(}Y_{0}\cup\bigcup_{i\in K}Y_{i}\bigr{)}$
$\leq 0.$
This follows from the $k$-alternating condition ($k$-Alt) for $\varphi_{2}$ and sets $Y_{i}=\pi_{2}((A\times S_{2})\cap X_{i})$.
In general, if $\varphi_{1}$ is a coverage function, then by Proposition 2.3, $\varphi_{1}=\sum_{\emptyset\neq A\subseteq S_{1}}c_{A}\cdot\varphi_{A}$ for some nonnegative coefficients $c_{A}\in\mathbb{R}_{+}$ for all nonempty $A\subseteq S_{1}$. Define the tensor product of $\varphi_{1}$ and $\varphi_{2}$ to be
$\varphi_{1}\otimes\varphi_{2}\coloneqq\sum_{\emptyset\neq A\subseteq S_{1}}c_{A}\cdot(\varphi_{A}\otimes\varphi_{2}).$
Then, the resulting function over $S_{1}\times S_{2}$ is a tensor product of $\varphi_{1}$ and $\varphi_{2}$. Indeed, $(\varphi_{1}\otimes\varphi_{2})(Y_{1}\times Y_{2})=\sum_{\emptyset\neq A\subseteq S_{1}}c_{A}\cdot(\varphi_{A}\otimes\varphi_{2})(Y_{1}\times Y_{2})=(\sum_{A\cap Y_{1}\neq\emptyset}c_{A})\cdot\varphi_{2}(Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$. Finally, observe that $\varphi_{1}\otimes\varphi_{2}$ is $k$-alternating as it is a nonnegative combination of $k$-alternating functions. ∎
As noted earlier, ($k$-Alt) for $k=1$ corresponds to the increasing property, while for $k=2$ it is equivalent to the increasing submodularity of the function. Thus, by applying Theorem 3.16 to the case when $k=2$, we get the following.
Corollary 3.17.
Any coverage function $\varphi_{1}$ and increasing submodular function $\varphi_{2}$ have an increasing submodular tensor product.
Remark 3.18.
Note that if we do not assume $\varphi_{2}$ to be increasing, then a submodular tensor product might not exist. As an example, let $S_{1}\coloneqq\{a,b\}$, $\varphi_{1}\coloneqq\varphi_{S_{1}}$, $S_{2}\coloneqq\{1,2\}$, and $\varphi_{2}\colon S_{2}\to\mathbb{R}_{+}$ defined by $\varphi_{2}(\emptyset)=\varphi_{2}(S_{2})=0$ and $\varphi_{2}(1)=\varphi_{2}(2)=1$, and assume that $\varphi$ is a submodular tensor product of $\varphi_{1}$ and $\varphi_{2}$. Then, using the notation $a_{i}=(a,i)$ and $b_{i}=(b,i)$ for $i\in\{1,2\}$, we have
$\varphi(\{a_{1},a_{2},b_{1}\})$ $=\varphi(\{a_{1},a_{2},b_{1}\})+\varphi(\{a_{1}\})-1\leq\varphi(\{a_{1},a_{2}\})+\varphi(\{a_{1},b_{1}\})-1=0+1-1=0,$
$\varphi(\{a_{1},a_{2},b_{1}\})$ $=\varphi(\{a_{1},a_{2},b_{1}\}+\varphi(\{b_{1},b_{2}\})\geq\varphi(\{b_{1}\})+\varphi(\{a_{1},a_{2},b_{1},b_{2}\})=1+0=1,$
a contradiction.
Another interesting special case is when both $\varphi_{1}$ and $\varphi_{2}$ are coverage functions.
Corollary 3.19.
Any two coverage functions have a coverage tensor product.
Remark 3.20.
In [7, Example 4.17], it is shown that the rank function of a matroid is a coverage function if and only if it is a partition matroid. Combining this observation with Theorem 3.16 implies that any matroid has a tensor product with any partition matroid, which is a well-known result on tensor products.
The smallest non-partition matroid is $U_{2,3}$, whose rank function is therefore not a coverage function. Interestingly, any matroid that is not a partition matroid contains $U_{2,3}$ as a minor. Additionally, recalling that any two matroids representable over the same field admit a tensor product explains why the smallest counterexample of a pair of matroids without a tensor product is the pair $U_{2,3}$ and the Vámos matroid.
3.5 Extension to the infinite case
The main results of Section 3.1 and Section 3.2 can be extended to pairs of bounded submodular set functions defined on infinite set algebras. Essentially the same ideas work as in the finite cases, i.e., in the proofs of 3.1 and 3.2, but the reasoning is slightly more technical, and we have to integrate instead of taking sums. Moreover, we will need to integrate with respect to nonnegative charges. We follow the definitions of [30]. Let $(J,\mathcal{B})$ be a set algebra, $\alpha\colon\mathcal{B}\to\mathbb{R}_{+}$ be a nonnegative charge, and $f\colon J\to\mathbb{R}_{+}$ be a nonnegative bounded measurable function. Define
$\hat{\alpha}(f)\coloneqq\int_{0}^{\infty}\alpha(\{f\geq t\})\,\mathop{}\!\mathrm{d}t.$
Note that $\hat{\alpha}$ is linear in the sense that $\hat{\alpha}(c_{1}f_{1}+c_{2}f_{2})=c_{1}\hat{\alpha}(f_{1})+c_{2}\hat{\alpha}(f_{2})$ for all nonnegative bounded measurable functions $f_{1},f_{2}$, and $c_{1},c_{2}\in\mathbb{R}_{+}$. Furthermore, $\hat{\alpha}$ is positive, meaning that $\hat{\alpha}(g)\geq\hat{\alpha}(f)$ if $g(x)\geq f(x)$ for all $x\in J$.
Remark 3.21.
In [30], the integral is defined for any set function $\alpha$ with bounded variation, and any bounded measurable function $f$.
Let $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$ be set algebras and denote their product space by $(J,\mathcal{B})$. Let $\varphi_{1}\colon\mathcal{B}_{1}\to\mathbb{R}_{+}$, $\varphi_{2}\colon\mathcal{B}_{2}\to\mathbb{R}_{+}$ be bounded submodular functions satisfying $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$, and let $\alpha_{1}\colon\mathcal{B}_{1}\to\mathbb{R}_{+}$, $\alpha_{2}\colon\mathcal{B}_{2}\to\mathbb{R}_{+}$ be charges. To make it easier to draw parallels with the finite case, by a slight abuse of notation, we denote by $\varphi_{2}(\pi_{2}(Z_{x}))$ the function $x\mapsto\varphi_{2}(\pi_{2}(Z_{x}))$ and by $\varphi_{1}(\pi_{1}(Z^{y}))$ the function $y\mapsto\varphi_{1}(\pi_{1}(Z^{y}))$ for any set $Z\in\mathcal{B}$. Any set $Z$ of the product algebra has the form $Z=\cup_{i=1}^{n}A_{i}\times B_{i}$, where $A_{i}\in\mathcal{B}_{1}$ and $B_{i}\in\mathcal{B}_{2}$ for $i\in[n]$. Consider the finite algebras $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ generated by the sets $\{A_{i}\}_{i=1}^{n}$ and the sets $\{B_{i}\}_{i=1}^{n}$, respectively. It is not difficult to see that $\varphi_{2}(\pi_{2}(Z_{x}))$ is constant on the atoms of $\mathcal{A}_{1}$ and $\varphi_{1}(\pi_{1}(Z^{y}))$ is constant on the atoms of $\mathcal{A}_{2}$. Consequently, both functions take only finitely many values, and the preimages of these values belong to $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively. As $\mathcal{A}_{i}\subseteq\mathcal{B}_{i}$ for $i\in[2]$, we conclude that both functions are measurable. Similarly to (2), we can define
$b(Z)\coloneqq\hat{\alpha}_{1}(\varphi_{2}(\pi_{2}(Z_{x})))+\hat{\alpha}_{2}(\varphi_{1}(\pi_{1}(Z^{y})))-\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ (10)
for all $Z\in\mathcal{B}$. While this definition does not look symmetric at first, note that $\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ and $\hat{\alpha}_{2}(\alpha_{1}(\pi_{1}(Z^{y})))$ agree on the semi-algebra of product sets and thus agree on the generated algebra, see e.g. [34, Theorem 3.5.1 (ii)]. Observe that if $Z=Y_{1}\times Y_{2}$ for some $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, then
$\hat{\alpha}_{1}(\varphi_{2}(\pi_{2}(Z_{x})))$ $=\alpha_{1}(Y_{1})\varphi_{2}(Y_{2}),$
$\hat{\alpha}_{2}(\varphi_{1}(\pi_{1}(Z^{y})))$ $=\varphi_{1}(Y_{1})\alpha_{2}(Y_{2}),$
$\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ $=\alpha_{1}(Y_{1})\alpha_{2}(Y_{2}).$
First we extend the statement of 3.1.
Theorem 3.22.
Let $\varphi_{1},\varphi_{2}$ be nonnegative submodular functions over the set algebras $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$, respectively, satisfying $\varphi_{1}(\emptyset)=0$ and $\varphi_{2}(\emptyset)=0$. Furthermore, let $\alpha_{1}\colon\mathcal{B}_{1}\to\mathbb{R}_{+}$ and $\alpha_{2}\colon\mathcal{B}_{2}\to\mathbb{R}_{+}$ be charges such that $\alpha_{1}(S_{1})=\varphi_{1}(S_{1})$ and $\alpha_{2}(S_{2})=\varphi_{2}(S_{2})$. Then, the function $b$ defined in (10) is a submodular coupling of $\varphi_{1}$ and $\varphi_{2}$.
Proof.
By the definition of $b$, we have
$b(Y_{1}\times Y_{2})$ $=\alpha_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\alpha_{2}(Y_{2})\cdot\varphi_{1}(Y_{1})-\alpha_{1}(Y_{1})\cdot\alpha_{2}(Y_{2})$
$=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})-(\varphi_{1}(Y_{1})-\alpha_{1}(Y_{1}))\cdot(\varphi_{2}(Y_{2})-\alpha_{2}(Y_{2}))$
$=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2}),$
if $Y_{1}=J_{1}$ or $Y_{2}=J_{2}$, hence $b$ is indeed a coupling. To prove submodularity, note that for fix $x\in J_{1}$ and $y\in J_{2}$, the functions $\varphi_{2}(\pi_{2}(Z_{x}))$ and $\varphi_{1}(\pi_{1}(Z^{y}))$ are submodular while $\alpha_{2}(\pi_{2}(Z_{x}))$ is modular, hence $b$ is submodular by the linearity and nonnegativity of $\alpha_{1}$ and $\alpha_{2}$. ∎
To extend the result of 3.2, we first have to define the measurable analogue of a base polytope. Let $(J,\mathcal{B})$ be a set algebra and $\varphi\colon\mathcal{B}\to\mathbb{R}_{+}$ be a polymatroid function. We say that a nonnegative charge $\alpha\colon\mathcal{B}\to\mathbb{R}_{+}$ is a minorizing charge if $\alpha(A)\leq\varphi(A)$ for all $A\in\mathcal{B}$. A minorizing charge is called basic if $\alpha(J)=\varphi(J)$. The set of all basic minorizing charges is denoted by $\operatorname{bmm}(\varphi)$. It was proved in [30, Corollary 5.3.] that $\operatorname{bmm}(\varphi)$ is nonempty.
Theorem 3.23.
Let $\varphi_{1}$, $\varphi_{2}$ be polymatroid functions on measurable spaces $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$, respectively. Then, there exists a polymatroid coupling $\varphi$ of $\varphi_{1}$ and $\varphi_{2}$ on the product space $(J,\mathcal{B})$.
Proof.
Choose $\alpha_{1}\in\operatorname{bmm}(\varphi_{1})$ and $\alpha_{2}\in\operatorname{bmm}(\varphi_{2})$, and define $b$ as in (10).
Claim 3.24.
We have $b(\emptyset)=0$ and $b(Y_{1}\times Y_{2})\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})$ for each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, where equality holds if $\varphi_{1}(Y_{1})=\alpha_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\alpha_{2}(Y_{2})$.
Proof.
The equality $b(\emptyset)=0$ follows from $\varphi_{1}(\emptyset)=\varphi_{2}(\emptyset)=0$. For each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$, $\alpha_{1}(Y_{1})\leq\varphi_{1}(Y_{1})$ and $\alpha_{2}(Y_{2})\leq\varphi_{2}(Y_{2})$, thus we get
$b(Y_{1}\times Y_{2})=\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})-(\varphi_{1}(Y_{1})-\alpha_{1}(Y_{1}))\cdot(\varphi_{2}(Y_{2})-\alpha_{2}(Y_{2}))\leq\varphi_{1}(Y_{1})\cdot\varphi_{2}(Y_{2}).$
Here equality holds if $\varphi_{1}(Y_{1})=\alpha_{1}(Y_{1})$ or $\varphi_{2}(Y_{2})=\alpha_{2}(Y_{2})$. ∎
The function $b$ is not necessarily monotone. Let $\varphi(Z)\coloneqq\inf_{Z^{\prime}\supseteq Z}b(Z^{\prime})$ for all $Z\in\mathcal{B}$.
Claim 3.25.
$\varphi$ is increasing and submodular.
Proof.
Clearly, $\varphi$ is increasing by definition. To prove submodularity, let $Z_{1},Z_{2}\in\mathcal{B}$. Then, for any $\varepsilon>0$ and for $i=1,2$, there exists $Z^{\prime}_{i}\supseteq Z_{i}$ such that $\varphi(Z_{i})+\varepsilon\geq b(Z^{\prime}_{i})$. We get
$\varphi(Z_{1})+\varphi(Z_{2})$ $\geq b(Z^{\prime}_{1})+b(Z^{\prime}_{2})+2\varepsilon$
$\geq b(Z^{\prime}_{1}\cap Z^{\prime}_{2})+b(Z^{\prime}_{1}\cup Z^{\prime}_{2})+2\varepsilon$
$\geq\varphi(Z_{1}\cap Z_{2})+\varphi(Z_{1}\cup Z_{2})+2\varepsilon.$
The claim follows by taking $\varepsilon\to 0$. ∎
Claim 3.26.
$\varphi(Y_{1}\times Y_{2})=b(Y_{1}\times Y_{2})$ holds for each $Y_{1}\in\mathcal{B}_{1}$ and $Y_{2}\in\mathcal{B}_{2}$.
Proof.
We need to show that $b(Z)\geq b(Y_{1}\times Y_{2})$ holds if $Y_{1}\times Y_{2}\subseteq Z\in\mathcal{B}$. Using that $\varphi_{2}$ is increasing and $\alpha_{2}(X)\leq\varphi_{2}(X)$ holds for any $X\in\mathcal{B}_{2}$, we get
$\hat{\alpha}_{1}(\varphi_{2}(\pi_{2}(Z_{x})))$ $=\hat{\alpha}_{1}(\mathbbm{1}_{Y_{1}}\cdot\varphi_{2}(\pi_{2}(Z_{x})))+\hat{\alpha}_{1}(\mathbbm{1}_{J_{1}\setminus Y_{1}}\cdot\varphi_{2}(\pi_{2}(Z_{x})))$
$\geq\alpha_{1}(Y_{1})\cdot\varphi_{2}(Y_{2})+\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}((Z\setminus(Y_{1}\times J_{2}))_{x}))).$
Similarly,
$\hat{\alpha}_{2}(\varphi_{1}(\pi_{1}(Z^{y})))$ $=\hat{\alpha}_{2}(\mathbbm{1}_{Y_{2}}\cdot\varphi_{1}(\pi_{1}(Z^{y})))+\hat{\alpha}_{2}(\mathbbm{1}_{J_{2}\setminus Y_{2}}\cdot\varphi_{1}(\pi_{1}(Z^{y})))$
$\geq\varphi_{1}(Y_{1})\cdot\alpha_{2}(Y_{2})+\hat{\alpha}_{2}(\mathbbm{1}_{J_{2}\setminus Y_{2}}\cdot\alpha_{1}(\pi_{1}(Z^{y})))$
$=\varphi_{1}(Y_{1})\cdot\alpha_{2}(Y_{2})+\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}((Z\setminus(J_{1}\times Y_{2}))_{x}))).$
Finally,
$\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}(Z_{x})))$ $=\alpha_{1}(Y_{1})\cdot\alpha_{2}(Y_{2})+\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}((Z\setminus(Y_{1}\times Y_{2}))_{x})))$
$\leq\alpha_{1}(Y_{1})\cdot\alpha_{2}(Y_{2})+\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}((Z\setminus(Y_{1}\times J_{2}))_{x})))+\hat{\alpha}_{1}(\alpha_{2}(\pi_{2}((Z\setminus(J_{1}\times Y_{2}))_{x}))).$
Combining these three inequalities the claim follows. ∎
Combining Claims 3.243.25 and 3.26, we deduce that $\varphi$ is a polymatroid coupling of $\varphi_{1}$ and $\varphi_{2}$. This concludes the proof of the theorem. ∎
4 Universal functions
In this section we show that, under certain conditions, if there exists a coupling of any two functions with some property $P$, then there exists a universal function with property $P$, i.e., a function that contains any finite function with property $P$ as a quotient.
4.1 Preparations
If $\mathcal{Q}=\{Q_{i}\}_{i=1}^{n}$ is a finite measurable partition of $J$, i.e., $\{Q_{i}\}_{i=1}^{n}$ is a partition and $Q_{i}\in\mathcal{B}$ for $i\in[n]$, then the quotient $\varphi/\mathcal{Q}\colon 2^{[n]}\to\mathbb{R}$ denotes the function defined by $(\varphi/\mathcal{Q})(X)=\varphi(\bigcup_{i\in X}Q_{i})$ for $X\subseteq[n]$. The notion of coupling can be naturally extended to more than two functions. Let $(J_{1},\mathcal{B}_{1}),\dots,(J_{n},\mathcal{B}_{n})$ be set algebras and denote their product space by $(J,\mathcal{B})$. Furthermore, let $\varphi_{i}\colon\mathcal{B}_{i}\to\mathbb{R}$ be set functions for $i\in[n]$. We say that a function $\varphi\colon\mathcal{B}\to\mathbb{R}$ is a coupling of $\varphi_{1},\dots,\varphi_{n}$ if $\varphi(S_{1}\times\dots\times S_{i-1}\times X_{i}\times S_{i+1}\times\dots\times S_{n})=\varphi_{i}(X_{i})\cdot\prod_{\ell\in[n]\setminus\{i\}}\varphi_{j}(J_{\ell})$ for every $X_{i}\in\mathcal{B}_{i}$, $i\in[n]$.
Let $(J_{1},\mathcal{B}_{1})$ and $(J_{2},\mathcal{B}_{2})$ be two set algebras. The functions $\varphi_{1}\colon\mathcal{B}_{1}\to\mathbb{R}$ and $\varphi_{2}\colon\mathcal{B}_{2}\to\mathbb{R}$ are isomorphic if there exists a bijection $f\colon J_{1}\to J_{2}$ such that $f$ and $f^{-1}$ are measurable and $\varphi_{2}(f(X))=\varphi_{1}(X)$ for all $X\in\mathcal{B}_{1}$. By a property of set functions, we mean an isomorphism invariant class of set functions. We call a property $P$ of set functions finitary if a function $\varphi\colon\mathcal{B}\to\mathbb{R}$ is in $P$ if and only if $\varphi/\mathcal{Q}$ is in $P$ for any finite measurable partition $\mathcal{Q}$ of $J$. A finitary property $P$ is closed if for any measurable space $(J,\mathcal{B})$ and any finite measurable partition of $J$, the set of functions $\{f\colon\mathcal{B}\to\mathbb{R}\mid f/\mathcal{Q}\in P\}$ is closed. We call a finitary property $P$ extendable if for any $\varphi\colon 2^{[n]}\to\mathbb{R}$ with property $P$ and any finite, pairwise disjoint sets $A_{1},\ldots,A_{n}$ with union $A$, there exists a function $\psi\colon 2^{A}\to\mathbb{R}$ with property $P$ that extends $\varphi$ in the sense that that $\varphi(S)=\psi(\bigcup_{i\in S}A_{i})$ for all $S\subseteq[n]$. We call a finitary property $P$ liftable if any two finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. Note that liftability implies that any finite number of finite functions with property $P$ admit a coupling with property $P$ on the product of their ground sets. We say that property $P$ is $b$-bounded for some $b\in\mathbb{R}_{+}$ if the range of any function with property $P$ is contained in $[-b,b]$. Finally, a normalized function with property $P$ is called universal for property $P$, if it contains all finite, normalized functions with property $P$ as a quotient.
We need the following well-known lemma (which follows from e.g. [25, Problem 4.43]) and include a proof for completeness.
Lemma 4.1.
There exists a family of subsets $\mathcal{H}\subseteq 2^{\mathbb{Z}}$ with $|\mathcal{H}|=\mathfrak{c}$ such that for any disjoint finite subsets $\mathcal{H}_{1},\mathcal{H}_{2}\subseteq\mathcal{H}$, we have
$\bigcap_{H\in\mathcal{H}_{1}}H\cap\bigcap_{H\in\mathcal{H}_{2}}H^{c}\neq\emptyset.$
Proof.
Let
$\mathcal{R}\coloneqq\bigl{\{}A\subseteq\mathbb{R}\mid A=\bigcup_{i=1}^{n}[p_{i},q_{i}]\text{ for some $n\in\mathbb{Z}_{+}$ and rational numbers $p_{i},q_{i}$ for $i\in[n]$}\bigr{\}}.$
Note that $\mathcal{R}$ is countable, hence it is enough to find an appropriate system with $\mathcal{H}\subseteq 2^{\mathcal{R}}$. For any $x\in\mathbb{R}$, let $\mathcal{H}_{x}\coloneqq\{A\in\mathcal{R}\mid x\in A\}$. For any disjoint finite subsets $X_{1}\subseteq\mathbb{R}$ and $X_{2}\subseteq\mathbb{R}$ we can clearly find $A\in\mathcal{R}$ with $X_{1}\subseteq A$ and $X_{2}\cap A=\emptyset$. Hence the set system $\{\mathcal{H}_{x}\}_{x\in\mathbb{R}}$ satisfies the condition of the lemma. ∎
The proof of 4.2 relies on a compactness argument. For a set algebra $(J,\mathcal{B})$, the space of set functions $\{f\colon\mathcal{B}\to\mathbb{R}\}$ can be identified with the product topological space $\prod_{A\in\mathcal{B}}\mathbb{R}$. For any $r\colon\mathcal{B}\to\mathbb{R}_{+}$, the space of functions
$\bigl{\{}f\colon\mathcal{B}\to\mathbb{R}\ \big{|}\ |f(A)|\leq r(A)\text{ for all }A\in\mathcal{B}\bigr{\}}=\prod_{A\in\mathcal{B}}[-r(A),r(A)]$
is compact, as the product of compact sets is compact by Tychonoff’s theorem [38]. We will use that a topological space $X$ is compact if and only if $\bigcap_{C\in\mathcal{C}}C\neq\emptyset$ for any collection $\mathcal{C}$ of closed subsets of $X$ such that $\bigcap_{C\in\mathcal{C}^{\prime}}C\neq\emptyset$ for all finite subcollections $\mathcal{C}^{\prime}\subseteq\mathcal{C}$.
4.2 Proof of existence
Now we are ready to state the main result of this section.
Theorem 4.2.
For any finitary, closed, extendable, liftable and $b$-bounded property $P$ for some $b\geq 1$, there exists a universal function on the set algebra $(\mathbb{Z},2^{\mathbb{Z}})$.
Proof.
Let $\mathcal{H}\subseteq 2^{\mathbb{Z}}$ be a set given by 4.1. Let $\Phi$ be a set of representatives of functions with property $P$ with finite domain whose size is a power of two, i.e., for all such isomorphism class $\Phi$ contains exactly one function. For a function $\varphi\in\Phi$, denote by $S_{\varphi}$ the domain of $\varphi$. We may assume that for any two distinct $\varphi,\psi\in\Phi$, their domains are disjoint, that is, $S_{\varphi}\cap S_{\psi}=\emptyset$. Since $P$ is extendable, it is enough to find a function $F\colon 2^{\mathbb{Z}}\to\mathbb{R}$ with property $P$ that contains every member of $\Phi$ as a quotient.
Up to isomorphism, there are continuum real-valued functions on finite domain, hence $|\Phi|\leq\mathfrak{c}$. Consequently, we can choose $H_{1}^{\varphi},\dots,H_{n}^{\varphi}\in\mathcal{H}$ for all $\varphi\in\Phi$, where $n=\log_{2}(|S_{\varphi}|)$, such that $H_{i}^{\varphi_{1}}=H_{j}^{\varphi_{2}}$ only if $\varphi_{1}=\varphi_{2}$ and $i=j$. Furthermore, for all $\varphi\in\Phi$ and all $a\in S_{\varphi}$, define $I_{a}\subseteq[n]$ and
$T_{a}\coloneqq\bigcap_{i\in I_{a}}H_{i}^{\varphi}\cap\bigcap_{i\in[n]\setminus I_{a}}\left(H_{i}^{\varphi}\right)^{c},$
where $n=\log_{2}(|S_{\varphi}|)$, such that $I_{a}\neq I_{b}$ for distinct $a,b\in S_{\varphi}$. It is enough to prove the existence of a function $F\colon 2^{\mathbb{Z}}\to\mathbb{R}$ such that $F$ has property $P$ (which is equivalent to every finite factor having property $P$) and
$F\bigl{(}\bigcup_{a\in A}T_{a}\bigr{)}=\varphi(A)$
for all $\varphi\in\Phi$ and $A\subseteq S_{\varphi}$, as the latter condition guarantees that it contains every finite function with property $P$ as a quotient. Note that, by the property of the family $\mathcal{H}$ given by 4.1, $\bigcup_{a\in A}T_{a}\neq\bigcup_{b\in B}T_{b}$ for any $\varphi,\psi\in\Phi$ and $A\subseteq S_{\varphi}$, $B\subseteq S_{\psi}$.
The function space $\{f\colon\mathcal{B}\to\mathbb{R}\mid|f(A)|\leq b\text{ for all $A\in 2^{\mathbb{Z}}$}\}$ is compact. The condition that $F/\mathcal{Q}$ has property $P$ is closed for every finite measurable partition $\mathcal{Q}$ of $\mathbb{Z}$. Furthermore, conditions of the form $F(\bigcup_{a\in A}T_{a})=\varphi(A)$ are also closed. Hence, it is enough to prove that for any finite number of these conditions, we can find a function satisfying all of them. Let $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{n}$ be finite partitions of $\mathbb{Z}$, and let $\bigcup_{a\in A_{i}}T_{a}$ be sets for some $A_{1},\ldots,A_{m}$, where there exist distinct $\varphi_{1},\ldots,\varphi_{k}\in\Phi$ such that $A_{i}\subseteq S_{\varphi_{f(i)}}$ for all $1\leq i\leq m$ and some function $f\colon[m]\to[k]$. We need to show that there exists a function $F^{\prime}\colon 2^{\mathbb{Z}}\to\mathbb{R}$ such that $F^{\prime}(\bigcup_{a\in A_{i}}T_{a})=\varphi_{f(i)}(A_{i})$ for $1\leq i\leq m$, and $F^{\prime}/\mathcal{Q}_{i}$ has property $P$ for $1\leq i\leq n$.
[…]
$\eta(A)=\bigcup_{(a_{i})_{i=1}^{k}\in A}\bigl{(}\bigcap_{i=1}^{k}T_{a_{i}}\bigr{)}$
for all $A\in 2^{S}$. Now we can define the function $F_{1}\colon\mathcal{A}_{1}\to\mathbb{R}$ by $F_{1}(A)=\varphi(\eta^{-1}(A))$ for all $A\in\mathcal{A}_{1}$. By definition, this satisfies all equalities of the form $F_{1}(\bigcup_{a\in A_{i}}T_{a})=\varphi_{f(i)}(A_{i})$, and $F_{1}$ has property $P$ as $\varphi$ has property $P$. Let $\mathcal{A}_{2}$ be the algebra generated by the sets of $\mathcal{A}_{1}$ and the sets appearing in one of the partitions $\mathcal{Q}_{i}$ for $1\leq i\leq n$. This is a refinement of $\mathcal{A}_{1}$ still having a finite number of atoms, hence, by extendability, there exists a function $F_{2}\colon\mathcal{A}_{2}\to\mathbb{R}$ with property $P$ that extends $F_{1}$, i.e., $F_{2}|_{\mathcal{A}_{1}}=F_{1}$. It follows that $F_{2}/\mathcal{Q}_{i}$ has property $P$ for $1\leq i\leq n$. Finally, an arbitrary extension $F^{\prime}$ of $F_{2}$ to the domain $2^{\mathbb{Z}}$ concludes the proof of the theorem. ∎
4.3 Applications
We show two applications of 4.2. First, we show that there exists an universal increasing submodular set function, using 3.2. Any increasing, submodular set function which has value 00 on the empty set (apart form the constant 0) can be normalized, hence such a universal function contains all increasing, submodular set functions with value 00 on the empty set up to a constant multiplier.
Corollary 4.3.
There exists a universal increasing submodular set function.
Proof.
Let $P$ be the property of being normalized, increasing and submodular. To apply 4.2, we have to show that $P$ is a finitary, closed, extendable, liftable, 1-bounded property. For a set function $\varphi\colon\mathcal{B}\to\mathbb{R}$ on $(J,\mathcal{B})$, having property $P$ means
$\varphi(A)\leq\varphi(B)\ \ \text{ and }\ \ \varphi(C\cap D)+\varphi(C\cup D)\leq\varphi(C)+\varphi(D)$ (11)
for all $A,B,C,D\in\mathcal{B}$ with $A\subseteq B$, and being normalized. Hence, it is clear that $P$ is finitary, and 1-boundedness also holds by monotonicity and $\varphi(J)=1$. Liftability follows from 3.2. Note that every inequality in (11) is satisfied by a closed subset of $\{f\colon\mathcal{B}\to\mathbb{R}\}$, and the functions with normalized property is also closed subset, hence the set of functions having property $P$ in $\{f\colon\mathcal{B}\to\mathbb{R}\}$ is closed, as it is the intersection of closed subsets. Consequently, $\{f\colon\mathcal{B}\to\mathbb{R}\mid f/\mathcal{Q}\in P\}$ is closed for all finite measurable partition, as it is a projection of $\{f\colon\mathcal{B}\to\mathbb{R}\}$ to the algebra generated by $\mathcal{Q}$, proving that $P$ is closed. Finally, if $\varphi\colon 2^{[n]}\to\mathbb{R}$ has property $P$, and $A_{1},A_{2},\ldots,A_{n}$ are finite, pairwise disjoint sets with union $A$, then the set function $\psi\colon 2^{A}\to\mathbb{R}$ defined by $\psi(X)\coloneqq\varphi(\nu(X))$ for all $X\subseteq A$, where $\nu(X)\coloneqq\{i\in[n]\mid X\cap A_{i}\neq\emptyset\}$ for all $X\subseteq A$, is an extension with property $P$. Indeed, $\psi$ is clearly an increasing extension, and for any $X,Y\subseteq A$,
$\psi(X)+\psi(Y)$ $=\varphi(\nu(X))+\varphi(\nu(Y))$
$\geq\varphi(\nu(X)\cap\nu(Y))+\varphi(\nu(X)\cup\nu(Y))$
$\geq\varphi(\nu(X\cap Y))+\varphi(\nu(X\cup Y))$
$=\psi(X\cap Y)+\psi(X\cup Y).$
Hence property $P$ is extendable as well, so we can apply 4.2 to finish the proof of the corollary. ∎
In particular, 4.3 implies that there exists a function $F\colon 2^{\mathbb{Z}}\to\mathbb{R}$ that contains all normalized matroid rank functions as a quotient.
As a second application, we show that there exists a universal function for coverage functions using 3.19. The proof relying on 4.2 is only existential; later, in Section 4.4, we describe an explicit construction.
Corollary 4.4.
There exists a universal coverage function.
Proof.
The proof is similar to that of 4.3. The coverage property is given by the inequalities of ($k$-Alt), hence this property is finitary, closed, and by monotonicity it is 1-bounded. 3.19 gives liftability. It remains to prove that if $\varphi\colon 2^{[n]}\to\mathbb{R}$ is a coverage function and $A_{1},A_{2},\ldots,A_{n}$ are finite, pairwise disjoint sets with union $A$, then we can extend $\varphi$ on $A$. By 2.3, we can write
$\varphi=\sum_{\begin{subarray}{c}\emptyset\neq X\subseteq[n]\end{subarray}}c_{X}\varphi_{X},$
where $\varphi_{X}$ is defined as in (1) and $c_{X}\geq 0$ for all $\emptyset\neq X\subseteq[n]$. Let $u(X)\coloneqq\bigcup_{i\in X}A_{i}$ for all $X\subseteq[n]$ and define
$\psi\coloneqq\sum_{\begin{subarray}{c}\emptyset\neq X\subseteq[n]\end{subarray}}c_{X}\varphi_{u(X)}.$
Then $\psi$ is a coverage function by 2.3, and it clearly extends $\varphi$, proving extendability. Hence the corollary follows by 4.2. ∎
Remark 4.5.
Note that the existence of a universal function for $b$-bounded submodular functions does not follow from our results, as 3.1 does not ensure that the coupling of $b$-bounded submodular functions remains $b$-bounded. Furthermore, even achieving extendability is nontrivial, as it is unclear how to extend a $b$-bounded submodular function while preserving the bound.
4.4 Construction for coverage functions
By Corollary 4.4, finite normalized coverage functions admit a universal function. However, the proof of Theorem 4.2 offers little insight into the form of such a function. In this section, we construct an explicit coverage function so that any finite coverage function can be obtained as its quotient.
Theorem 4.6.
Let $\mathcal{B}$ be the Borel sigma-algebra on $\mathbb{R}$ and $\lambda$ be the Lebesgue measure. For any $A\in\mathcal{B}$, define $\Phi(A)=\lambda(\{\operatorname{frac}(a)\mid a\in A\})$. Then, $\Phi$ is a universal coverage function.
Proof.
Note that for any $A\in\mathcal{B}$, we have $\{\operatorname{frac}(a)\mid a\in A\}=\bigcup_{k\in\mathbb{Z}}\{a-k\mid a\in A\cap[k,k+1)\}$ which is a measurable set; hence, $\Phi(A)$ is well-defined. First we show that $\Phi$ is indeed a coverage function. To see this, let $A_{0},A_{1},\dots,A_{k}\in\mathcal{B}$ be arbitrary. Define $\hat{A_{i}}\coloneqq\{\operatorname{frac}(a)\mid a\in A_{i}\}$ for $i=0,1,\dots,k$. Then, ($k$-Alt) for $\Phi$ and sets $A_{0},A_{1},\dots,A_{k}$ is equivalent to ($k$-Alt) for $\lambda$ and sets $\hat{A_{0}},\hat{A_{1}},\dots,\hat{A_{k}}$. Since any measure is a coverage function, the same holds for $\Phi$.
Now let $q$ be a positive integer and $\psi\colon 2^{[q]}\to\mathbb{R}$ be a normalized coverage function, that is, $\psi(\emptyset)=0$, $\psi([q])=1$ and $\psi$ satisfies ($k$-Alt) for all positive integer $k$ and choices of $A_{0},A_{1},\dots,A_{k}\in 2^{[q]}$. Since $\psi$ is increasing and normalized, we have $0\leq\psi\leq 1$. For ease of discussion, we use the notation $f(A)\coloneqq\{\operatorname{frac}(a)\mid a\in A\}$. Split the range $[0,1)$ into $q$ intervals $L_{1},\dots,L_{q}$ where $L_{i}=[\psi([i-1]),\psi([i]))$ for $i\in[q]$. Then, for each $i\in[q]$, the set $\operatorname{frac}^{-1}(L_{i})$ consists of infinitely many intervals $\{J_{i}^{m}\}_{m=1}^{\infty}$ where, for ease of discussion, we define $J_{i}^{m}=\emptyset$ for $m\in\mathbb{Z}_{>0}$ if $\psi([i-1])=\psi([i])$.
We construct a measurable partition $\mathcal{Q}=\{Q_{1},\dots,Q_{q}\}$ of $\mathbb{R}$ for which $\psi=\Phi/\mathcal{Q}$ in $q$ successive steps. In step $t$, we construct a function $\phi_{t}$ over $2^{[q]}$ satisfying the following:
  1. (A)
    $\phi_{t}(I)=\psi(I)$ if $I\subseteq[t]$, $t\in I$,
  2. (B)
    $\phi_{t-1}(I)=\phi_{t}(I)$ if $I\subseteq[t-1]$.
Note that (A) and (B) together implies $\phi_{q}=\psi$. Indeed, let $I\subseteq[q]$ be an arbitrary set. If $\max I=q$, then $\phi_{q}(I)=\psi(I)$ holds by (A). Otherwise, we have $\phi_{q}(I)=\phi_{\max I}(I)=\psi(I)$, where the first equality follows by (B) while the second follows by (A).
In Step 1, we add set $Q_{i}$ to be $J_{i}^{1}$ for each $i\in[q]$ and define $\phi_{1}(I)\coloneqq\Phi(\bigcup_{i\in I}Q_{i})$ for $I\subseteq[q]$. The function $\phi_{1}$ thus obtained is modular and satisfies $\phi_{1}\leq\psi$; in particular, if $\psi$ is modular then we have $\phi_{1}=\psi$. By construction, (A) is satisfied as $\phi_{1}(\{1\})=\Phi(Q_{1})=\Phi(J_{1}^{1})=\psi(\{1\})$. Even more, $\phi_{1}([t])=\psi([t])$ holds for every $t\in[q]$.
For $t\geq 2$, in Step $t$ we increase $Q_{t}$ by adding certain subsets $B_{t}^{I}\subseteq J^{t}_{\min I}$ for $\emptyset\neq I\subseteq[t-1]$, and then we define $\phi_{t}(I)\coloneqq\Phi(\bigcup_{i\in I}Q_{i})$ for $I\subseteq[q]$. This guarantees that $\phi_{t-1}(I)=\phi_{t}(I)$ if $I\subseteq[t-1]$, that is, (B) is satisfied. However, for (A) to be satisfied as well, the sets $B_{t}^{I}$ must be chosen carefully.
For $\emptyset\neq I\subseteq[t-1]$, let us define
$b_{t}^{I}=\sum_{J\subseteq I\cup\{t\}}(-1)^{|J|+1}\psi\big{(}([t-1]\setminus I)\cup J\big{)}.$
The role of these quantities is that eventually we will set $B_{t}^{I}$ such that $b_{t}^{I}=\Phi(B_{t}^{I})$, but first, we prove two technical claims providing lower and upper bounds on $b_{t}^{I}$.
Claim 4.7.
$b_{t}^{I}\geq 0$.
Proof.
The statement follows by ($k$-Alt) for $\psi$ and $k\coloneqq|I|+1$ with the choice $A_{0}=[t-1]\setminus I$ and $A_{1},\dots,A_{k}$ being the singletons of $I\cup\{t\}$. ∎
Claim 4.8.
$b_{t}^{I}\leq b_{\max{I}}^{I\setminus\{\max I\}}-(b_{\max{I}+1}^{I}+...+b_{t-1}^{I})$.
Proof.
By slight rearrangement and the definition of $b_{t}^{I}$, we need to show that
$\sum_{J\subseteq I}(-1)^{|J|+1}\psi\big{(}([\max I]\setminus I)\cup J\big{)}\geq\sum_{i=\max I+1}^{t}\sum_{J\subseteq I\cup\{i\}}(-1)^{|J|+1}\psi\big{(}([i-1]\setminus I)\cup J\big{)}.$
This can be vastly simplified, as the right hand side is a telescoping sum. Indeed, if for a given choice of $i=i_{0}$ we have a term ±$\pm\psi(([i_{0}-1]\setminus I)\cup J)$ with $i_{0}\in J$, then for $i=i_{0}+1$ we encounter the same term if we drop $i_{0}$ from $J$, with an opposite sign. This yields that all terms cancel except for the ones with $i=\max I+1$, $i\notin J$, and $i=t$, $i\in J$. However, the first source of terms precisely corresponds to the left hand side, hence that cancels as well. Thus the inequality simplifies to
$0\geq\sum_{J\subseteq I}(-1)^{|J|}\psi\big{(}([t]\setminus I)\cup J\big{)}.$
The latter inequality follows from ($k$-Alt) for $\psi$ and $k\coloneqq|I|$ with the choice $A_{0}=[t]\setminus I$ and $A_{1},\dots,A_{k}$ being the singletons of $I$. ∎
With the help of Claims 4.7 and 4.8, we show that there exist subsets $B_{t}^{I}\subseteq J_{\min I}^{t}$ for $\emptyset\neq I\subseteq[t-1]$ satisfying $b_{t}^{I}=\Phi(B_{t}^{I})$ and being located so that the growth $\phi_{t}(I)-\phi_{t-1}(I)$ for $I\subseteq[t]$, $t\in I$ is precisely what is needed for (A), as will be shown in Claim 4.10. For a set $X\subseteq[0,1)$, we denote by $X^{c}$ the complement of $X$, that is, $X^{c}=[0,1)\setminus X$.
Claim 4.9.
For $\emptyset\neq I\subseteq[t-1]$, there exists $B_{t}^{I}\subseteq\operatorname{frac}^{-1}((\bigcap_{i\in I}f(Q_{i}))\cap(\bigcap_{i\in[t-1]\setminus I}f(Q_{i})^{c}))\cap J^{t}_{\min I}$ such that $\Phi(B_{t}^{I})=b_{t}^{I}$.
Proof.
First observe that since $\Phi(B_{t}^{I})=\lambda(f(B_{t}^{I}))\geq 0$, we need $b_{t}^{I}\geq 0$, which holds by Claim 4.7. On the other hand, we need to show that there is enough unallocated space in $J_{\min I}^{t}$ for the set $B_{t}^{I}$. That is, since for the desired $B_{t}^{I}$ we have $f(B_{t}^{I})\subseteq(\bigcap_{i\in I}f(Q_{i}))\cap(\bigcap_{i\in[t-1]\setminus I}f(Q_{i})^{c})$, we need
$b_{t}^{I}\leq\lambda\Big{(}\big{(}\bigcap_{i\in I}f(Q_{i})\big{)}\cap\big{(}\bigcap_{i\in[t-1]\setminus I}f(Q_{i})^{c}\big{)}\Big{)}.$
By definition, we have
$b_{\max{I}}^{I\setminus\{\max I\}}=\lambda\Big{(}\big{(}\bigcap_{i\in I}f(Q_{i})\big{)}\cap\big{(}\bigcap_{i\in[\max I]\setminus I}f(Q_{i})^{c}\big{)}\Big{)},$
and the sets $B_{j}^{I}$ for $j=\max I+1,\dots,t-1$ have pairwise disjoint images in $f(B_{\max{I}}^{I\setminus\{\max I\}})$ of measure $b_{j}^{I}$. Thus, the sufficient and necessary upper bound on $b_{t}^{I}$ can be rewritten as
$b_{t}^{I}\leq b_{\max{I}}^{I\setminus\{\max I\}}-(b_{\max{I}+1}^{I}+...+b_{t-1}^{I}),$
which follows by Claim 4.8. ∎
Let $B_{t}^{I}$ be sets provided by Claim 4.9. We define $B_{t}\coloneqq\bigcup_{\emptyset\neq I\subseteq[t-1]}B_{t}^{I}$ and update $Q_{t}$ by adding $B_{t}$ to it.
Claim 4.10.
For $I\subseteq[t-1]$, we have $\phi_{t}([t]\setminus I)=\psi([t]\setminus I)$.
Proof.
We can express
$\Phi(\bigcup_{i\in[t]\setminus I}Q_{i})=\Phi(\bigcup_{i\in[t-1]\setminus I}Q_{i})+\Phi(Q_{t})+\sum_{\emptyset\neq J\subseteq I}\Phi(B_{t}^{J}).$
As $\Phi(Q_{t})=\psi([t])-\psi([t-1])$ by definition, $\phi_{t-1}([t-1]\setminus I)=\psi([t-1]\setminus I)$ by (A) and (B), and $\Phi(B_{t}^{J})=b_{t}^{J}$, we get that $\phi_{t}([t]\setminus I)=\psi([t]\setminus I)$ holds if and only if
$\psi([t]\setminus I)=\psi([t-1]\setminus I)+\psi([t])-\psi([t-1])+\sum_{\emptyset\neq J\subseteq I}b_{t}^{J}.$
For $I=\emptyset$, the equation is satisfied due to Step 1. Otherwise, using the definition of $b_{t}^{I}$, one can count the occurrences of $\psi([t-1]\setminus I^{\prime})$ and $\psi([t]\setminus I^{\prime})$ for each $\emptyset\subseteq I^{\prime}\subseteq I$ on both sides. As coefficients, we get sums of the form $\sum_{i=0}^{|I\setminus I^{\prime}|}(-1)^{i}\binom{|I\setminus I^{\prime}|}{i}$ which is 00, except for the cases when $I^{\prime}=\emptyset$ or $I^{\prime}=I$. These exceptions yield the remainder terms $\psi([t]\setminus I)$, $\psi([t-1]\setminus I)$, $\psi([t])$, $\psi([t-1])$ with appropriate signs. ∎
By Claim 4.10, property (A) is also satisfied by $\phi_{t}$. Therefore, after $q$ steps, the above procedure provides a subpartition $Q_{1},\dots,Q_{q}$ of $\mathbb{R}$ satisfying $\psi(I)=\phi_{q}(I)=\Phi(\bigcup_{i\in I}Q_{i})$ for $I\subseteq[q]$. This subpartition, however, can be extended to a proper partition of $\mathbb{R}$ by adding $\operatorname{frac}^{-1}(L_{i})\setminus(\bigcup_{i=1}^{q}Q_{i})$ to $Q_{i}$ for $i\in[q]$. This does not change the function $\phi_{q}$, concluding the proof of the theorem. ∎
5 Conclusions and open problems
In this paper, we introduce the notion of coupling for set functions, providing a novel operation that combines two set functions into a single one. The proposed framework fills the gap arising from the fact that matroids do not always have a tensor product, and provide a construction that is applicable in every case while still possessing properties similar to the tensor product. As applications, we gave new necessary conditions for the representability of matroids, and established the existence of universal functions from which any submodular function on a finite ground set can be obtained by taking quotients.
We close the paper by mentioning some open problems:
  1. 1.
    Theorem 3.2 shows that two $2$-alternating functions always admit a $2$-alternating coupling, while Corollary 3.19 provides an analogous result for coverage functions. This leads to a natural question: Do two $k$-alternating functions always have a $k$-alternating coupling?
  2. 2.
    We have seen that the common information property is a necessary condition for linearity, which in turn implies Ingleton’s inequality. Additionally, we showed that having a tensor product with the uniform matroid $U_{2,3}$ is also a necessary condition for linearity, and it too implies Ingleton’s inequality. It remains an intriguing open problem to clarify how these two necessary conditions are related.
  3. 3.
    The proofs of Theorems 3.1 and 3.2 provided constructions for coupling two submodular functions and two polymatroid functions, respectively. However, it is not difficult to find examples showing that not all couplings can be obtained through these constructions. The description of the convex hull of couplings remains an interesting open problem.
Acknowledgment.
The authors are grateful to Miklós Abért, Márton Borbényi, Máté András Pálfy and László Márton Tóth for helpful discussions. András Imolay was supported by the Rényi Doctoral Fellowship of the Rényi Institute and by the EKÖP-24 University Excellence Scholarship Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant numbers ADVANCED 150556 and ELTE TKP 2021-NKTA-62, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115. This work was supported in part by EPSRC grant EP/X030989/1.
Data availability
No data are associated with this article. Data sharing is not applicable to this article.